A Bibliography of Publications of Gilbert Wright ‘Pete’ Stewart III

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@ieee.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.22

Abstract

This bibliography records publications of Gilbert Wright ‘Pete’ Stewart III.

Title word cross-reference

AX + XB = C [BS72]. Ax = λBx [MS71, Ste72, Ste75b]. B [Ste10b]. CS [Ste82]. LU [Ste93f, Ste97d]. n [Ste16a]. QR [CPS97a, CPS97b, DGKS76, MS93a, MS93b, Ste77b, Ste84a, Ste87d, Ste92a, Ste93f, Ste94d, Ste95c, Ste98d, Ste99b, Ste10a]. ULV [Ste93h]. URV [AGS91a, AGS91b, GS91, GBS92, LOSW93, LOSW94, Ste91e, Ste94i]. UT V [Ste94a].

-algorithm [Ste87d]. -Arnoldi [Ste10b].

1 [LS95]. 10-16 [BF90]. 13-17 [IEE02]. 15th [GW94]. 18th [BS86]. 1967 [DH69]. 1974 [Mil75]. 1975 [BR76]. 1976 [HW76]. 1977 [Ric77]. 1978 [DS79, dBG78, Ple81]. 1982 [KR83]. 1983 [ICH84]. 1985 [All86]. 1986 [BS86, IP87]. 1988 [ER88]. 1991 [IEE91]. 1992 [MGD93]. 1993 [BCEP94, GW94].

2002 [IEE02]. 2nd [Ste95g].

384 [Ste70a, Ste70e].

432 [BS72].

506 [Ste76a].

776 [BS97].

844 [BPS05]. 879 [LS09]. 8th [ER88].

92 [IEE92].

= [GS95].

A. [Ste79d]. Academy [Mil75]. Accelerating [Ste69a]. Accept [Ste67b]. Accuracy [Ste97a]. ACM [BS72, Ste70a]. Acoustics [IEE91, IEE92, IEE02]. adaptive [LOSW93]. Addendum [Ste02a]. Adjugate [Ste98b]. Adjustable [Ste09a]. Adjusting [Ste02b]. Advanced [MGD93, Kin99, Ste98f]. Advances [SGW^{+}89, dBG78, GGL94, vH97]. after [Ste96]. Afternotes [Ste96, Ste98f, Kin97, Kin99, Ser97]. Aggregated [Ste83a]. Alamos [TTB^{+}88]. Algebra [DH69, HS87, MGD93, Ste88b, Bir80, MP93, Ste99a]. Algorithm [BS72, BS79, GOS15, HC03, MS93a, MS93b, MS73, MS78, Ste70a, Ste70b, Ste71a, Ste79a, Ste85a, Ste01b, Ste02a, Ste05, MS71, MS72, SSM89, SMC08, Ste87d, Ste90g, Ste91e, Ste92c, BS97, BPS05, FW82, LS09, Ste70e, Ste76a]. Algorithms [DGKS76, Ste86b, Ste98d, Ste01c, BCcC^{+}88, OS85, Ste98g, Ste99b, Vac91, Ste97a]. Alston [Ste93a]. Ambiguity [SSS89]. AMS [BF90]. AMS-IMS-SIAM [BF90]. analogue [CJ04]. Analyses [CPS96, CPS97a, CPS97b]. Analysis [IP87, SL68, Ste68, Ste69e, Ste69f, Ste71a, Ste92a, Ste05, Ste11b, dBG78, Bir80, BF90, GW94, JS01, Mil75, Ste96, Ste98f, Vac91, Kin97, Ste93b, Kin99, Ser97]. Analysts [Ste97b]. analytic [Ste71c]. Anecdotes [TTB^{+}88]. Application [Ste80b, TTH^{+}84]. Applications [MP93, MGD93, BF90, OS98, Vac91]. Applied [Bir80, MP93]. Approach [Ste00b]. approaches [AMPS83]. Approximate [Ste71b, Ste02c]. Approximating [HC03, JS01].

Approximation [GHS87, GHS88, Ste99c]. Approximations [Ste67b, Ste98d, BPS05, Ste99b, Ste01a]. April [HW76, IP87]. architectures [BCcC^{+}88]. Argonne [BR76]. Arithmetic [Ste09a]. Army [BF90]. Arnoldi [Ste00a, Ste09b, Ste10b]. ARPACK [Ste00a]. Arrays [OS87]. arrival [AGS91a, AGS91b]. Arrowhead [OS90]. Art [IP87]. Artificial [AMPS83]. Asilomar [Che91]. Aspects [DH69]. Assignment [OS86]. Associated [Ste73b]. Asymptotic [Ste84a]. August [MGD93, Ros74].

B [Bir80, Cam81, Par81]. Backward [Ste69e, Ste02c]. Band [RS98]. Bartels [SMC08]. based [AMPS83, LOSW93]. Bases [Ste88a, Ste12]. Basic [Ste98g]. Basis [GMS^{+}85]. Bau [Ste99a]. Bauer [FGH^{+}06]. be [Ste81a]. Behavior [Ste84a, Ste80a]. Behavioural [SMC08]. Belgium [MGD93]. Below [TTB^{+}88]. Bibliographical [Ste76b]. Bigradients [HS69b]. Biography [Duf10]. Birmingham [IP87]. Bit [TTB^{+}88]. Block [BS79, MS93a, MS93b, Ste08, vMS94, Ste95e, vS96]. Book [Bir80, Fra95, Kin97, Kin99, Par81, Ser97, Ste78a, Ste93b, Ste97a, Ste99a, Ste00a]. Bounds [O’L90, Ste71b, Ste73b, Ste77b, Ste79c, Ste83a, Ste91g, Ste98e, Ste02c]. Broad [RS98]. Broad-Band [RS98]. Bug [TTB^{+}88]. Building [Ste03a, OPSW86]. Bunch [Bir80, Cam81, Par81]. Byte [TTB^{+}88].

C [Bir80, Cam81, Par81]. Calculate [BS92, BS97, Ste78c]. Calculating [BS74, FW82, Ste75d, Ste76a]. calculation [Ste90b]. calculations [Ste81b]. calculator [Bir80]. calculators [Ste81b]. California [Wou89, Che91, IEE92]. Camille [Ste16a]. Canada [IEE91]. cancellation [LHS13]. Canonical [Ste16b]. capital [Ste87c]. Carl [Fra95]. Carolina [BCEP94]. Case [Ste11c].

Centenary [BCEP94]. Center [Ric77, dBG78]. Centre [IEE91]. Certain [Ste73b]. Chains [Ste83a, Ste84c, Ste90i, Ste90c, SZ91b, Ste93d, Ste94e, SSM94, Ste97c, MP93, Ste93g, Ste95g, Ste95g]. Charles [Ste87e, Ste93b]. Cholesky [CPS96, Ste79a, Ste93f, Ste97d]. Classifier [HBC^{+}00]. Closed [Ste71b]. Coefficients [Ste77d]. College [Ste96]. Collinearity [Ste86a, Ste87a, Ste87b]. Collins [BS86]. Colorado [BS86]. column [Ste84b, Ste90b]. combination [Fra95, GS95]. combinationis [Fra95, GS95]. commentaries [KO10]. commentary [Ste11a]. Comments [HS69a]. Communication [Ste86b, Ste90a]. Compact [Bir80]. companion [Ste71c]. Compilers [Ste76f]. complement [Ste95d]. Completely [SSM94]. Complex [Ste85b, Ste86c, Pri04]. COMPSTAT [ER88]. Computable [Ste83a]. Computation [OSvdG89, Ste98d, Ste00b, CH90, OS85, Ste99b]. computational [ER88, Bir80]. Computations [BR76, OSvdG86, Ste73d, Ste76f, Ste90a, Ste95g, Ste87e]. Computer [All86, HW76, SGW^{+}89, TTH^{+}84, ICH84, All86, BS86]. Computers [Che91, Bir80, OPSW86]. Computing [ACM98, BS79, GOS15, HS87, OS90, SS74, Ste82, Ste85a, Ste88b, TTB^{+}88, GSV86b, Ste76e, Ste83b, BPS05]. Concerning [Ste88a]. Condition [CMSW79, Ste80b, SZ91a, Ste90b]. conditioned [Ste81a]. Conducted [Ric77, dBG78, DH69]. Conference [BCEP94, Che91, GW94, IEE91, IEE92, IP87, KR83, Mil75, BF90, IEE02, Che91]. Congress [Ros74]. Conjugate [Ste73a, Ste75a]. Constrained [Ste81a]. constraints [Ste97e]. Constructive [DH69]. Continuity [Ste69b]. Convergence [HC03, Ste70d, Ste74a, Ste75a, Ste94f, Ste98a, Ste09b, OS98]. Copenhagen [ER88]. Cornelius [BCEP94].

Correcting [Ste76c]. Corrigendum [Ste71d, Ste86c]. Crab [OPSW86]. Criterion [Ste09b]. CS [Ste16b]. Curiosity [Ste88a]. Curriculum [FGH^{+}06].

Data [OS85]. Data-flow [OS85]. David [Ste99a]. Davidon [Ste67b]. Debugging [TTB^{+}88]. December [BCEP94]. Decomposable [SSM94]. Decomposing [MS78]. Decomposition [GOS15, MS93a, MS93b, Ste85a, Ste91f, Ste93e, Ste93h, Ste99c, AGS91a, AGS91b, GS91, GBS92, HC05, LOSW93, Ste82, Ste83b, Ste90e]. Decompositional [Ste00b]. Decompositions [Ste84a, Ste94a, Ste94d, Ste97f, Ste98c, Ste94i, Ste95c, Ste98g]. Definite [OSV79, Ste79c, Ste81a]. Deflation [Ste81c]. Degeneracy [GKS76b, GKS76a, Ste76c, Ste84d]. Derivatives [MS88, Ste67b]. derived [Ste03b]. Detecting [Ste76c]. detection [LHS13]. Determinacy [OS87]. determinants [HS69b]. Determining [Ste93c]. Developing [BDP^{+}98a, BDP^{+}98b]. Development [Cow84, Ste77c, TTH^{+}84]. Diagonalization [BS79]. Diagonally [BS74]. Difference [Ste67b]. digital [MS79]. Dimensions [Ste16a]. Direct [SZ91b]. Direction [AGS91a, AGS91b, Ste73a]. Direction-of-arrival [AGS91a, AGS91b]. disks [Ste67a]. Division [Ste71a, Ste85b, Ste86c, Pri04]. Domain [RS98]. Dominant [BS92, BS97, BS74, Ste78c]. DOMINO [OSvdG86, OSvdG89]. Dongarra [Bir80, Cam81, Par81]. Downdates [Ste94g, Ste95f]. Downdating [Ste79a, Ste10a]. Duff [Ple81]. Dundee [GW94]. Dynamic [LHS13, SMC08].

Early [Ste93e]. Eckart [GHS87, GHS88]. Economical [Ste76d]. edited [Ple81]. Editor [BCcC^{+}88]. Effects [Ste77d, Ste79a].

Efficient [GOS15, Pri04, Ste80b, Ste98d, MS79, Ste99b]. Eigenproblems [Ste01b, Ste02a, LS09, OS98]. eigenspaces [JS01, Ste12]. Eigensystems [Ste01c, Ste01d]. EIGENTEST [LS09]. Eigenvalue [MS73, OSV79, Ste72, Ste73b, Ste74c, Ste75b, Ste76b, Ste78b, Ste79c, SZ91a, Ste00a, MS71, MS72]. Eigenvalues [FW82, OS90, Ste70a, Ste70e, Ste76a, SZ91a, Ste91g, Ste98e, Ste04]. Eigenvectors [BS74, MS88, OS90, Ste70a, Ste70e, Ste75d, Ste69a]. elementary [Ste96]. Elements [Ste74b]. Elimination [Ste74b, Ste93d, Ste94b, Ste95a, Ste97f, Ste98c]. Elsner [Ste04]. Elsner-like [Ste04]. Enough [Ste11c]. Environment [OSvdG86]. equality [Ste97e]. Equation [BS72, Ste92b]. Equations [GS76, Ste81c, Ste11a, Ste73a]. Error [Ste69e, Ste71a, Ste71b, Ste73b, Ste79a, Ste83a, Ste86a, Ste92a, Ste93c, Ste05, BF90, Fra95, GS95, Ste80a, Ste02c]. erroribus [Fra95, GS95]. Errors [HS72, Ste77d, Ste97b, vMS94, Ste90d, vH97, vS96]. errors-in-variables [vH97]. ESPRIT [LOSW93, LOSW94]. Essay [Ste16a]. Estimate [CMSW79]. Estimating [OSV79]. estimation [AGS91a, AGS91b]. Estimators [Ste80b]. Example [Ste86b]. EXCHNG [FW82, Ste76a]. Excitement [HS72]. Expansion [SS93, Ste84e]. Exponential [Ste92a]. Extraction [RS98]. Extreme [Ste75a].

F [Ste87e]. F2 [FW82, Ste70a, Ste70e, Ste76a]. F4 [BS72]. fact [Ste96]. factoring [HS69a, Ste69c, Ste70c, Ste73c]. Factorization [CPS96, CPS97a, CPS97b, DGKS76, HS71, OS86, Ste77b, Ste79a, MS79, Ste91e]. Factorizations [Ste93f, Ste10a]. Factors [Ste97d]. Fan [Ste67a]. Fashioned [Ste03a]. February [ACM98, GGL94]. fifth [Che91]. File [Ste91a, Ste91c]. Finding [Ste69d, Ste70d, Ste71d, Ste80a]. First [TTB^{+}88, TTH^{+}84]. Flap [Ste09a]. Floating [Ste09a, LHS13]. Floating-Point [Ste09a, LHS13]. Florida [IEE02]. flow [OS85]. Formulation [Ste69e]. Fort [BS86]. FORTRAN [BS92, BS97, Ste78c, FW82, Ste76a]. Foundation [BF90]. Four [Ste99b]. Fox [MS78]. Francisco [IEE92]. FREDHOLM [Ste11a]. Friedrich [Fra95, FGH^{+}06]. FTP [Ste91a]. FTPing [Ste91b]. function [Bir80]. functions [Ste71c]. Fundamental [DH69, Ste11a].

G [Bir80, Cam81, Fra95, Kin97, Kin99, LS95, Par81, Ple81, Ser97]. G. [Duf10, KO10, Ste11a]. Gauss [Fra95, Ste94b, Ste95a]. Gaussian [Ste74b, Ste93d, Ste94b, Ste95a, Ste97f, Ste98c]. Gene [Ste87e]. General [Ste91c]. Generalization [GHS87, GHS88, Ste67a, Ste73c, Ste01a]. Generalized [GOS15, MS73, SS74, Ste69b, Ste75b, Ste76b, Ste78b, Ste79c, Ste92b, MS71, MS72, Ste83b, Ste04, Ste06]. Generation [Ste80b]. generator [LS09]. Geometry [Ste16a]. Gershgorin [Ste67a, Ste75b]. given [MP93]. goes [Kin99, Ste98f]. Golub [Ste87e]. Graded [SZ91a, Ste94d, Ste95c, Ste01d]. Gradients [Ste75a]. graduate [Kin99, Ste98f]. Gram [DGKS76, Ste05, Ste08]. groups [SSS89]. Grove [Che91]. Guide [DBMS79, Ste00a, Bir80, Cam81, Par81].

H [Ste87e]. hand [Ste81b]. hand-held [Ste81b]. Hankel [HS69b]. harmonic [CJ04]. Harvard [HW76]. Havsbad [KR83]. heat [SL68]. Heinz [Ste78a]. Held [KR83, BF90, DS79, ER88, GGL94, IP87, MP93, Ste81b, Wou89]. Henrici [Bir80]. Hermitian [Ste69a, Ste76e, Ste85a, Ste91g, Ste98e]. Hessenberg [FW82, Ste76a, Ste95e, vMS94, vS96]. Hessians [Ste81a]. hierarchical [AMPS83]. High [ACM98]. High-Performance [ACM98]. Higham [Ste97a]. HILBERT [Ste11a]. History [Ste93e]. Horn [Ste93b]. Hotel [All86, IEE91]. Householder [MS78, Ste79d, Ste93a]. HP [Bir80]. HQR3 [FW82, Ste76a]. Hyatt [DS79]. Hyperbolic [SS98]. hypergeneralized [Ste06].

Iain [Ple81]. IBM [DH69]. ICASSP [IEE91, IEE92]. ICASSP-91 [IEE91]. ICASSP-92 [IEE92]. Identification [HS72]. IEEE [IEE02]. IFIP [Ros74]. II [Mil75, Vac91, Ste73c, Ste01c]. III [Ric77, Ste99a]. IMA [GGL94, IP87]. IMA/SIAM [IP87]. image [MS79]. Impacts [SGW^{+}89]. implementation [Ste87d]. Implementing [SSM89]. Implicit [Ste81c]. Implicitly [Ste00a]. Impossible [TTB^{+}88]. IMS [BF90]. Incorporating [Ste70b]. Incremental [Ste90b]. Independent [Ste77d, Ste80a]. Industrial [Bir80]. Inequalities [Ste95b, Ste90h]. Information [Ros74]. Input [HS72]. Institute [HW76, MP93, MGD93]. Integers [Ste88a]. Integral [Ste11a]. intelligence [AMPS83]. Interface [All86, BS86, HW76]. interior [SSM89]. International [BCEP94, IEE91, IEE92, ICH84, Ste95g, IEE02]. Introduction [Ste73d]. invariance [Ste84b]. Invariant [BS92, BS97, Ste71b, Ste78c, Ste87c, Ste76e]. Invention [TTB^{+}88]. Inverse [SS74, Ste69b, Ste75c]. Inverses [O’L90, Ste77a, Ste89]. Irish [Mil75]. Isolated [Ste75a]. Italy [ICH84]. Iteration [Ste75d, SSM94, Ste69a, Ste70c, Ste76e]. Iterations [Ste74a, Ste94f, HS69a, Ste69c, Ste73c]. Iterative [KS99, SS74, Ste95b, GGL94, Ste90h].

J [Bir80, Cam81, Par81, Ste97a]. Jacobi [Ste85a]. Jacobi-Like [Ste85a]. January [MP93, Wou89]. Java [ACM98, BDP^{+}98a, BDP^{+}98b]. Jeep [Ste91c]. John [Bir80]. Johnson [Ste93b]. Joint [IP87, BF90]. Jonathan [TTB^{+}88]. Jordan [Ste16a]. July [GW94]. June [BF90, DH69, GW94].

Kentucky [All86]. Kernels [RS98].

Knoxville [DS79]. Krylov [Ste02a, Ste01b, Ste02c].

Laboratory [BR76, DH69]. Lanczos [BCEP94, Ste91d, Ste94c, Ste02b]. Large [MGD93, Ste74c, Ste76b, Ste00a, Ste01b, Ste02a, LS09, OS98, Ste90a]. Large-Scale [Ste00a, LS09]. Largest [OSV79]. Latin [Fra95]. Laurent [SS93]. leaks [Ste03b]. Learning [HBC^{+}00]. Least [GKS76b, GKS76a, Ste77a, Ste87a, Ste87b, Fra95, GS95, Ste75c, Ste90d, Ste97e, vH97]. lectures [Kin99, MP93, Ste96, Ste98f]. Lehmer [Ste71d, Ste69d]. Letter [BCcC^{+}88]. Leuven [MGD93]. Lexington [All86]. libraries [BDP^{+}98a, BDP^{+}98b]. Like [Ste85a, Ste04]. Linear [DS84, HS87, MP93, Ste71b, Ste77a, Ste77d, Ste81c, Ste88b, Ste95b, Bir80, Ste73a, Ste75c, Ste90h, Ste91d, Ste94c, Ste97e, Ste99a, MP93, MGD93]. LINPACK [Bir80, DBMS79, DS84, Ste77c, Cam81, Par81]. Lloyd [Ste99a]. Loan [Ste87e]. local [Ste12]. low [HC05]. low-rank [HC05]. LP [SSM89]. Ludwig [FGH^{+}06].

M [LS95]. M/G/1 [LS95]. Madison [Ric77, dBG78]. March [ACM98, All86, BS86, GGL94, IEE92, KR83, Ric77]. Markov [MP93, Ste95g, Ste83a, Ste84c, Ste90i, Ste90c, SZ91b, Ste93d, Ste93g, Ste94e, SSM94, Ste97c]. Marriott [IEE92]. Maryland [Ste96, OPSW86]. Massachusetts [HW76]. Mathematical [Cow84, Ric77, ICH84, Ric77].

Mathematics [Bir80, MP93, Ric77, dBG78]. Mathematik [Ste78a]. Matlab [Ste09a]. Matrices [BS74, MSS84, OS90, Ste70b, Ste75d, Ste80b, SZ91a, Ste94d, Ste97f, Ste98c, Ste98e, BPS05, HC05, Ste76e, Ste95c, Ste01d]. Matrix [BS92, BS97, BS72, BR76, CMSW79, DS79, FW82, GHS87, MS73, OSV79, OS86, Ste70a, Ste70e, Ste73d, Ste76a, Ste76f, Ste77b, Ste78c, Ste81b, Ste85a, SS90, Ste91g, Ste93b, Ste94h, Ste98b, Ste98g, Ste98d, Ste00b, Ste01c, LS09, MS71, MS72, MS79, OS85, Ple81, Ste69a, Ste82, Ste90a, Ste95d, Ste99b, Ste03c, KR83, Ste87e]. May [IEE91, IEE02, dBG78]. measurement [BF90]. medium [Wou89]. medium-scale [Wou89]. Memory [Ste76f, Ste03b]. Message [OSvdG86, Ste90a]. Method [GS76, SS74, Ste67b, Ste69d, Ste70d, Ste71d, Ste75a, SZ91b, Ste95b, Ste09b, Ste10b, JS01, OS98, Ste73c, Ste83b, Ste90h, Ste97e]. Methods [LS95, Ste75d, Ste02b, Bir80, GGL94, Ste73a, Ste00a]. minimis [Fra95]. Minimization [Ste67b, Bir80]. Minimum [GBS92]. Minimum-norm [GBS92]. minimus [GS95]. Minneapolis [GGL94]. Minnesota [GGL94]. MINRES [KS99]. Mirsky [GHS87, GHS88]. modeling [vH97]. modelling [SMC08]. models [BF90, MP93]. Modification [Ste67b]. Modifying [Ste74b]. module [OPSW86]. MOESP [SMC08]. Moler [Bir80, Cam81, Par81]. Moore [SS74]. Multiple [Ste74a, SZ91a]. multiplicity [Ste80a]. Multipoint [Ste74a, Ste94f]. multiprocessors [Wou89]. MUSIC [GS91].

N [Ste99a]. NASA [TTH^{+}84]. Nash [Bir80]. National [BF90, BR76]. NATO [MGD93]. Nearly [MSS84, Ste81c, Ste84c, Ste90c, SZ91b, SSM94, Ste93g]. Network [ACM98]. networks [SMC08]. neural [SMC08]. Nicholas [Ste97a]. Ninth [HW76]. Non [Ste85a, Ste76e]. Non-Hermitian [Ste85a, Ste76e]. Noninteger [Ste88a]. Nonlinear [GS76]. Nonsymmetric [BS92, BS97]. norm [GBS92]. North [BCEP94]. Note [Ste79b, Ste85b, Ste86c, Ste92b, Ste06]. November [Che91, DS79]. Null [GMS^{+}85, GOS15, Ste84b]. Number [CMSW79]. Numbers [SZ91a]. Numerical [HS87, HS71, IP87, Kin97, LS95, SS74, SL68, Ste68, Ste69f, Ste74c, Ste90i, Ste95g, Ste97a, Ste97b, Ste99a, Ste11b, dBG78, Bir80, BDP^{+}98a, BDP^{+}98b, CH90, Kin99, Mil75, Ser97, Ste96, Ste98f, GW94]. numerische [Ste78a].

Oblique [Ste11b, Ste11c]. obnoxiae

[GS95, Fra95]. observations [Fra95, GS95]. observationum [Fra95, GS95]. Office [BF90]. Old [Ste03a]. Old-Fashioned [Ste03a]. one [GS95]. Ontario [IEE91]. Operating [OSvdG89]. Operations [SGW^{+}89]. operator [Ste71c]. Operators [Ste71b]. optimization [AMPS83]. Order [Ste84e]. Ordering [FW82, Ste76a]. Origin [SS93, Ste70b, TTB^{+}88]. Orlando [IEE02]. Orthogonal [Ste80b, Ste69a]. Orthogonality [ES93]. Orthogonalization [Ste08]. orthonormal [Ste82].

Pacific [Che91]. Package [DS84, Ste09a]. Padé [HS69b]. Pair [Ste16b]. paper [Ste87c]. Papers [Ste11a, GGL94]. Parallel [BCcC^{+}88, GSV86b, OS86, OSvdG86, OSvdG89, Ste86b, Ste88b, OS85, OPSW86, Ste87d, Ste94i, Wou89]. Parametric [HS72]. Park [Ste96]. pars [GS95]. part [GS95, MP93]. participatory [Ste87c]. partitioned [Ste82]. Passing [OSvdG86, Ste90a]. Pencils [KR83, SS93, Ste94h]. Penrose [SS74]. Performance [ACM98, ICH84]. Periodic [HS72]. Perturbation [CPS96, CPS97a, CPS97b, Ste73b, Ste77b, Ste77a, Ste78b, Ste79b, Ste79c, Ste84e, SS90, Ste90d, Ste90e, Ste90f, Ste91f, Ste93d, Ste93f, Ste94h, Ste97d, LS00, Ste75c, Ste91e, Ste93g, Ste95d, Ste04]. Perturbations [MS88, Ste03c]. perturbed [Ste84b, Ste12]. Peter [Bir80]. Philadelphia [Bir80]. Phillips [RS98]. Pisa [ICH84]. Pite [KR83]. Pivot [Ste74b]. Pivoted [Ste98d, HC05, Ste99b]. Pivoting [SS98]. Plane [Ste76d]. plate [SL68]. Plaza [All86]. pocket [Bir80]. Point [Ste09a, LHS13, SSM89]. Points [Ste75a]. Polynomial [HS71, Ste69d, Ste70d, Ste71d, Ste71a, HS69a, Ste69c, Ste73c]. polynomials [Ste70c]. Positive [OSV79, Ste95d]. posterior [GS95]. powers [Ste03c]. pp [Bir80]. pp. [Bir80]. Precision [Ste09a]. Preface [GSV86a, Ste79d]. Presence [Ste93c, Ste80a]. presented [Ste96]. prior [GS95]. Problem [Ste72, Ste75b, Ste76b, Ste78b, Ste79c, MS71, SL68, Ste75c]. Problems [GKS76b, GKS76a, GGL94, MS73, Ste73b, Ste74c, Ste77a, Ste00a, AMPS83, MS72, Ste97e]. Proceedings [All86, BS86, BR76, DS79, IP87, KR83, MGD93, Ric77, Ros74, Wou89, dBG78, BF90, DH69, ER88, GW94, IEE02, ICH84, Mil75, Ple81, Ste95g, BCEP94, HW76]. Processing [IEE91, IEE92, IEE02, MS79, Ros74, Vac91, Wou89]. Products [Ste94d, Ste95c]. Program [Ste91a, TTB^{+}88]. Projection [MS78]. Projections [O’L90, Ste77a, Ste89]. Projectors [Ste11b, Ste06]. Properties [GMS^{+}85, SS74]. Pseudo [O’L90, Ste77a, Ste89]. Pseudo-Inverses [O’L90, Ste77a, Ste89]. punishment [Ste87c]. Purpose [Ste91c].

QLP [HC03, HC05, Ste99c]. QR [Ste70b]. Quasi [Ste05]. Quasi-Gram [Ste05]. queueing [MP93]. Queues [LS95]. Quotient [Ste02b, OS98].

R [Bir80, Cam81, Par81, Ste93b]. Radisson [All86]. Raleigh [BCEP94]. Random [Ste80b]. Rank [GKS76b, GKS76a, Ste76c, Ste84d, Ste93c, Ste93h, AGS91a, AGS91b, BPS05, GS91, GBS92, HC05, Ste91e]. Rank-Revealing [Ste93h, AGS91a, AGS91b, GS91, GBS92, Ste91e]. Rayleigh [JS01, MSS84, OS98, Ste01a, Ste02b]. Real [FW82, MGD93, Ste70a, Ste70e, Ste76a, Ste78c]. Real-Time [MGD93]. record [Che91]. recorded [Ste96]. Rectangular [Ste94h]. reduced [BPS05]. reduced-rank [BPS05]. Reducing [BS79]. Refinement [MSS84]. refining [Ste91e]. Regency [DS79]. Regression [Ste77d, Ste87a, Ste87b]. Regularization [KS99, RS98]. Rejoinder [Ste87b]. Related [Ste97f, Ste98c]. relative [LS00]. reliability [ICH84]. Reliable [CH90]. Remark [FW82, Ste70e]. Reminiscences [TTB^{+}88]. Renaissance [IEE02]. Reorthogonalization [DGKS76]. report [BCcC^{+}88]. Reports [Ste91b]. Representation [GMS^{+}85, Ste88a, Ste16b]. Requirements [HS72]. Research [DH69, Ric77, SGW^{+}89, Ste77c, dBG78, BF90]. Residual [Ste91g, Ste98e]. Resort [IEE02]. Restarted [Ste00a]. results [CJ04]. Revealing [Ste93h, AGS91a, AGS91b, GS91, GBS92, Ste91e]. Review [Bir80, Cam81, Fra95, Kin97, Par81, Ple81, Ste78a, Ste87e, Ste93b, Ste97a, Ste00a]. Reviews [Kin99, Ser97, Ste99a]. revisited [Ste03b]. Ritz [CJ04, JS01, MSS84, Ste01a]. Roger [Ste93b]. root [GS91, Ste80a]. root-finding [Ste80a]. root-MUSIC [GS91]. Rotations [Ste76d]. Rounding [Ste69e, Ste79a, Ste86a, vMS94, vS96]. Royal [Mil75]. Rüschlikon [DH69]. Rutishauser [Ste78a].

S [Ple81, Ste79d]. Saad [CJ04, Ste01a]. Samelson [Ste70c]. San [IEE92]. Scale [MGD93, Ste00a, LS09, Wou89]. Scaled [O’L90, Ste84a, Ste89]. Scaling [ES93, Ste86a, Pri04, Ste84b]. Scheduling [OS86]. scheme [Ste80a]. SCHMIDT [Ste11a, DGKS76, Ste05, Ste08]. school [Kin99, Ste98f]. Schur [Ste02a, Ste85a, Ste95d, Ste01b]. Science [BF90, HW76, All86, BS86]. Sciences [All86]. Scott [Ste93a]. Sebastião [Ste70d]. Secant [GS76, Ste73c]. Second [Ste84e]. selected [KO10]. selection [Ste90b]. Semidefinite [Ste10b, Ste95d]. Semiorthogonal [Ste02b]. Sensitivity [Ste72, Ste77d, Ste90c]. September [BR76, ICH84]. Sequential [Ste94g, Ste95f]. series [Ste96]. Set [Ste69e]. Seventeenth [All86]. Sheraton [IEE91]. Shifting [Ste71a]. Shifts [Ste70b]. SIAM [BF90, IP87]. Signal [IEE91, IEE92, IEE02, Vac91]. Signals [Che91, LOSW94]. Silva [Ste70d]. Simple [Ste91g, Ste98e]. Simultaneous [Ste75d, Ste76e]. Singular [MS93a, MS93b, SS93, Ste79b, Ste81c, Ste84a, Ste84e, Ste91f, Ste93e, Ste99c, LS00, Ste83b, Ste90e, Ste91e]. Sluggish [Ste94e, Ste97c]. Small [Ste84e]. Smooth [Ste12]. Society [Bir80]. Software [Cow84, Ric77]. Solution [BS72, SZ91b, Ste95g, Ste00a, SL68, Ste95e]. Solutions [Ste90i]. Solver [Ste03a]. Solving [DS84, GS76, SSM94, Ste95b, vMS94, AMPS83, Ste73a, Ste90h, vS96]. Some [HS69a, Ste68, Ste69c, Ste69f, Ste73c, MP93]. Sons [Bir80]. Sources [Cow84, RS98]. Space [GMS^{+}85, GOS15]. Spacecraft [TTH^{+}84]. Sparse [BR76, GGL94, Ste76b, Ste98d, Ste03a, BPS05, Ste99b, BR76, DS79, Ple81]. special [GSV86b]. Spectrum [Ste75a]. Speech [IEE91, IEE92, IEE02]. Squares [GKS76b, GKS76a, Ste77a, Ste87a, Ste87b, Ste75c, Ste90d, Ste97e, vH97]. SRRIT [BS92, BS97, Ste78c]. Stability [Ste94g, Ste97a, SS98, Ste95f]. Stable [DGKS76, GS76]. Stage [SSM94]. Standard [TTH^{+}84]. Stanford [Wou89]. State [IP87]. states [Ste93g]. Statistical [HS87, Ste88b, BF90]. Statistics [HW76, Ste94b, Ste95a, All86, BS86, ER88]. Stewart [Bir80, Par81, Cam81, CJ04, Duf10, Fra95, HC03, HC05, KO10, Kin97, Kin99, Ple81, SMC08, Ser97, Ste11a]. Stochastic [MSS84, Ste90f]. Stockholm [Ros74]. Storage [Ste76d]. Structure [Ste84c]. Structured [GGL94]. Study [MGD93]. Style [Ste91c]. subject [Fra95, GS95]. Sublinear [Ste98a]. Subroutine [BS97, Ste78c, BS92]. Subroutines [FW82, Ste76a]. Subspace [BS92, BS97, Ste90g, Ste92c]. Subspaces [BS79, Ste71b, Ste73b, Ste78c, Ste16b, LS00, Ste76e, Ste87c, Ste02c]. summer [BF90]. supercomputer [SSM89]. supplement [GS95]. supplementum [GS95]. support [BF90]. SVD [HC03, Vac91]. Sweden [KR83, Ros74]. Swift [TTB^{+}88]. Switzerland [DH69]. Sylvester [Ste92b]. Symmetric [OS90, Ste70a, Ste70b, Ste70e]. Symposium [All86, BS86, BR76, DS79, HW76, Ric77, dBG78, DH69, ER88, dBG78]. Synthetic [Ste71a]. Systaltic [OS87]. System [HBC^{+}00, OSvdG89]. Systems [Che91, DS84, HS72, Ste81c, vMS94, Ste73a, Ste90a, Ste91d, Ste94c, Ste95e, vS96].

table [HS69b]. Technical [Ste91b].

Technique [MSS84]. techniques [vH97]. Technology [HW76]. tend [Ste81a]. Tennessee [DS79]. Term [TTB^{+}88]. test [LS09]. testability [SSS89]. Theorem [DH69, GHS87, GHS88, LS00, Ste67a, Ste75c, Ste91e, Ste01a, Ste04]. Theoretic [Ste69e]. Theoria [GS95, Fra95]. Theory [GS95, Ste75b, Ste78b, SS90, Ste90f, Ste91f, Ste93d, Ste94h, Fra95, Ste90d, Ste90e]. thin [SL68]. Three [Ste11a]. Tikhonov [RS98]. Time [MGD93, RS98, LOSW94]. Time-Domain [RS98]. time-varying [LOSW94]. Toeplitz [RS98]. Topics [Ste68, Ste69f, Mil75, Ste93b]. Toronto [IEE91]. total [vH97]. Tour [Ste76b]. Towers [IEE91]. tracking [LOSW94, Ste90g, Ste92c]. Transfer [Ste91a, SL68]. transient [Ste93g]. Transients [Ste94e, Ste97c]. Translated [Ste11a]. Transportable [OSvdG89]. Treatment [Ste74c]. Trefethen [Ste99a]. Triangular [RS98, Ste97f, Ste98c]. Triangularization [SS98]. Tridiagonal [Ste70b]. Truncated [Ste98d, Ste99b]. Twenty [Che91]. Twenty-fifth [Che91]. Twice [Ste11c]. Two [Ste91g, SSM94, Ste98d, Ste98e, GS95]. Two-Stage [SSM94]. Type [LS95]. types [Ste03b].

U.S [BF90]. Uncoupled [MSS84, Ste84c, Ste90c, SZ91b]. University [HW76, IP87, Ric77, dBG78, Ste96]. Unreliable [Ste09b]. Updates [Ste94g, Ste95f]. Updating [DGKS76, GS91, Ste92a, Ste93h, Ste94i, GBS92, Ste90g, Ste92c]. Upper [FW82, Ste76a]. USA [IEE02]. User [Bir80, Cam81, DBMS79, Par81, Ste00a]. using [AGS91a, AGS91b, SMC08].

Value [MS93a, MS93b, Ste84a, Ste91f, Ste93e, Ste99c, Ste83b, Ste90e]. Values [Ste79b, Ste84e, Ste91e]. Variables [Ste77d, Ste97b, Ste90d, vH97]. Variant [GS76]. varying [LOSW94]. vectors [CJ04, Ste84b]. Virtual [Ste76f]. vitae [FGH^{+}06]. Volume [Ste98g, GSV86b, Ste01c]. Vorlesungen [Ste78a].

W [Bir80, Cam81, Duf10, Fra95, KO10, Kin97, Kin99, Par81, Ple81, Ser97, Ste11a]. weighting [Ste97e]. well [Ste81a]. Whence [TTB^{+}88]. Wiley [Bir80]. Windowing [Ste92a]. Wis [dBG78]. Wisconsin [Ric77, dBG78]. Word [TTB^{+}88]. works [KO10]. Workshop [ACM98, GGL94, MP93, Ste95g, BCcC^{+}88, ICH84, Wou89].

x [Bir80].

Year [MP93]. York [Bir80]. Young

[GHS87, GHS88].

Zero [Ste70d]. Zeros [Ste69d, Ste71d, Ste71a, Ste74a]. Zürich [DH69]. Zürich-Rüschlikon [DH69].

References

ACM:1998:AWJ

[ACM98] ACM, editor. ACM 1998 Workshop on Java for High-Performance Network Computing, February 28 and March 1, 1998. ACM Press, New York, NY, USA, 1998. ISBN ???? LCCN

                                                  ????  URL http://www.cs.
                                                  ucsb.edu/conferences/java98/

program.html. Possibly unpublished, except electronically.

Adams:1991:DAEa

[AGS91a] G. Adams, M. F. Griffin, and G. W. Stewart. Direction-of-arrival estimation using the rankrevealing URV decomposition. Technical Report CS-TR 2640, Department of Computer Science, University of Maryland, College Park, MD, USA, 1991. ???? pp.

Adams:1991:DAEb

[AGS91b] G. Adams, M. F. Griffin, and G. W. Stewart. Directionof-arrival estimation using the rank-revealing URV decomposition. In IEEE [IEE91], pages 1385–1388 (vol. 2). ISBN 0-7803-0003-3 (paperback), 0-7803-0004-1 (hardcover), 0-7803-0005-X (microfiche). LCCN TA365 I58 1991; TK7882.S65 I16 1991. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=150681.

Allen:1986:CSS

[All86] David M. Allen, editor. Computer science and statistics: Proceedings of the Seventeenth Symposium on the Interface of Computer Sciences and Statistics, Radisson Plaza Hotel, Lexington, Kentucky, March 17–19, 1985. North-Holland Publishing Co., Amsterdam, The Netherlands, 1986. ISBN 0-444-70018-8 (paperback). LCCN QA276.4 .S95 1985. URL http://

catalog.hathitrust.org/api/
volumes/oclc/13455449.html;
http://www.dtic.mil/dtic/
tr/fulltext/u2/a166019.pdf;
http://zbmath.org/?q=an:0647.
62008.

Anandalingam:1983:AIB

[AMPS83] G. Anandalingam, R. Mathieu, L. Pittard, and N. Sinha. Artificial intelligence based approaches for solving hierarchical optimization problems. In Impacts of Recent Computer Advances on Operations Research, pages 289– 301. North-Holland Publishing Co., Amsterdam, The Netherlands, 1983.

Buell:1988:LEP

[BCcC^{+}88] D. A. Buell, D. A. Carlson, Yuan chieh Chow, K. Culik, N. Deo, R. Finkel, E. N. Houstis, E. M. Jacobson, Z. M. Kedem, J. S. Kowalik, P. J. Kuekes, J. L. Martin, G. A. Michael, N. S. Ostlund, J. Potter, D. K. Pradhan, M. J. Quinn, G. W. Stewart, Q. F. Stout, L. Watson, and J. Webb. Letter to the Editor: Parallel algorithms and architectures: report of a workshop. The Journal of Supercomputing, 1(3):301–325, April 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0920-8542&
volume=1&issue=3&spage=301.

Brown:1994:PCL

[BCEP94] J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors. Proceedings of the Cornelius Lanczos International Centenary Conference, Raleigh, North Carolina, December 12–17, 1993, volume 73 of Proceedings in Applied Mathematics. SIAM (Society for Industrial and Applied Mathematics), Philadelphia, PA, USA, 1994. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Boisvert:1998:DNLa

[BDP^{+}98a] Ronald F. Boisvert, Jack J. Dongarra, Roldan Pozo, Karin A. Remington, and G. W. Stewart. Developing numerical libraries in Java. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.cs.ucsb.edu/
conferences/java98/papers/

jnt.pdf;   http://www.cs.ucsb.
edu/conferences/java98/papers/
jnt.ps;   http://www.netlib.
org/utk/people/JackDongarra/
PAPERS/jnt.ps;   http://www.
netlib.org/utk/people/JackDongarra/

pdf/jnt.pdf. Possibly unpublished, except electronically.

Boisvert:1998:DNLb

[BDP^{+}98b] Ronald F. Boisvert, Jack J. Dongarra, Roldan Pozo, Karin A. Remington, and G. W. Stewart. Developing numerical libraries in Java. Concurrency: Practice and Experience, 10(11– 13):1117–1129, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (elec-

tronic).  URL http://www3.
interscience.wiley.com/cgi-
bin/abstract?ID=10050395;
http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=10050395&
PLACEBO=IE.pdf. Special Issue:

Java for High-performance Network Computing.

Brown:1990:SAM

[BF90] Philip J. Brown and Wayne A.

Fuller, editors. Statistical analysis of measurement error models and applications: proceedings of the AMS-IMS-SIAM joint summer research conference held June 10-16, 1989, with support from the National Science Foundation and the U.S. Army Research Office, volume 112 of Contemporary Mathematics. American Mathematical Society, Providence, RI, USA, 1990. ISBN 0-8218-5117-9, 0-8218-7700-3 (ebook). ISSN 0271-4132 (print), 1098-3627 (electronic). LCCN QA275 .A47 1989. URL http:/

                                        /www.ams.org/conm/112/.

Birkhoff:1980:BRC

[Bir80] Garrett Birkhoff. Book review: Computational analysis with the HP 25 pocket calculator, by Peter Henrici, John Wiley, New York, 1977, 280 pp.; Compact numerical methods for computers: linear algebra and function minimization, by J. C. Nash, John Wiley & Sons, New York, 1979, x + 227 pp.; LINPACK: User’s guide, by J. J. Dongarra, J. R. Bunch, C. B. Moler, and G. W. Stewart, Society for Industrial and Applied Mathematics, Philadelphia, 1979 368 pp. Bulletin of the American Mathematical Society, 2(3):503–505, May 1980. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (elec-

                                        tronic). URL https://www.ams.
                                        org/journals/bull/1980-02-
                                        03/S0273-0979-1980-14785-0/
                                        S0273-0979-1980-14785-0.pdf.

Berry:2005:ACS

[BPS05] Michael W. Berry, Shakhina A. Pulatova, and G. W. Stewart. Algorithm 844: Computing sparse reduced-rank approximations to sparse matrices. ACM Transactions on Mathematical Software, 31(2):252–269, June 2005. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bunch:1976:SMC

[BR76] James R. Bunch and Donald J.

Rose, editors. Sparse Matrix Computations: Proceedings of the Symposium on Sparse Matrix Computations at Argonne National Laboratory on September 9–11, 1975. Academic Press Inc., New York, NY, USA, 1976. ISBN 0-12-141050-1. LCCN QA188 .S989 1975.

Bartels:1972:AAS

[BS72] Richard H. Bartels and G. W.

Stewart. ACM Algorithm 432: Solution of the matrix equation AX + XB = C [F4]. Communications of the Association for Computing Machinery, 15(9): 820–826, September 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Blevins:1974:CED

[BS74] M. M. Blevins and G. W. Stewart. Calculating the eigenvectors of diagonally dominant matrices. Journal of the Association for Computing Machinery, 21(2): 261–271, April 1974. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bavely:1979:ACR

[BS79] Connice A. Bavely and G. W.

Stewart. An algorithm for computing reducing subspaces by block diagonalization. SIAM Journal on Numerical Analysis, 16(2):359–367, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Boardman:1986:CSS

[BS86] Thomas J. Boardman and Irene M. Stefanski, editors. Computer science and statistics. Proceedings of the 18th Symposium on the Interface, Fort Collins, Colorado, March 1986. American Statistical Association, Washington, DC, USA, 1986. LCCN QA 276.4 S95 1986; TL695 N67

1992. URL http://zbmath.org/
?q=an:0638.62003.

Bai:1992:SFS

[BS92] Zhaojun Bai and G. W. Stewart. SRRIT — a FORTRAN subroutine to calculate the dominant invariant subspace of a nonsymmetric matrix. Technical Report CS-TR-2908, University of Maryland, College Park, MD, USA, May 1992. 25 pp. Also issued as University of Maryland Institute for Advanced Computer Studies UMIACS report TR-92-61.

Bai:1997:ASF

[BS97] Z. Bai and G. W. Stewart. Algorithm 776. SRRIT — a FOR-TRAN subroutine to calculate the dominant invariant subspace of a nonsymmetric matrix. ACM Transactions on Mathematical Software, 23(4):494–513, December 1997. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://

www.acm.org/pubs/citations/
journals/toms/1997-23-4/p494-
bai/.

Campbell:1981:RBU

[Cam81] Stephen L. Campbell. Review of Linpack User’s Guide by J. J. Dongarra, C. B. Moler, J. R. Bunch, and G. W. Stewart. Linear Algebra and Its Applications, 39(??):293–294, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379581903116.

Cox:1990:RNC

[CH90] M. G. Cox and S. Hammarling, editors. Reliable numerical computation. Oxford University Press, Oxford, UK, 1990. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th– 10th July 1987.

Chen:1991:CRT

[Che91] Ray R. Chen, editor. Conference record of the Twentyfifth Asilomar Conference on Signals, Systems and Computers: November 4–6, 1991, Pacific Grove, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-2470-1 (paperback), 0-8186-2471-X (microfiche), 0-8186-2472-8 (hardcover). LCCN TK 5102.5 A78 1991. URL

http://ieeexplore.ieee.org/
servlet/opac?punumber=382.

IEEE catalog number 91CH31120. IEEE Computer Society Press order number 2470.

Chen:2004:ARS

[CJ04] Guizhi Chen and Zhongxiao Jia.

An analogue of the results of Saad and Stewart for harmonic Ritz vectors. Journal of Computational and Applied Mathematics, 167(2):493–498, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042703009403.

Cline:1979:ECN

[CMSW79] A. K. Cline, C. B. Moler, G. W. Stewart, and J. H. Wilkinson. An estimate for the condition number of a matrix. SIAM Journal on Numerical Analysis, 16 (2):368–375, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (elec-

                                        tronic). URL http://zbmath.
                                        org/?q=an:0403.65012.

Cowell:1984:SDM

[Cow84] Wayne R. Cowell, editor. Sources and Development of Mathematical Software. Prentice-Hall Series in Computational Mathematics, Cleve Moler, Advisor. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1984. ISBN 0-13-823501-5. xii + 404 pp. LCCN QA76.95 .S68 1984.

Chang:1996:NPA

[CPS96] Xiao-Wen Chang, Christopher C. Paige, and G. W. Stewart. New perturbation analyses for the Cholesky factorization. IMA Journal of Numerical Analysis, 16 (4):457–484, October 1996. CO-DEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

                                        URL http://www3.oup.co.uk/

imanum/hdb/Volume_16/Issue_
04/160457.sgm.abs.html.

Chang:1997:PAF

[CPS97a] Xiao-Wen Chang, Christopher C.

Paige, and G. W. Stewart. Perturbation analyses for the QR factorization. SIAM Journal on Matrix Analysis and Applications, 18(3):775–791, July 1997. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29772.

Chang:1997:PAQ

[CPS97b] Xiao-Wen Chang, Christopher C.

Paige, and G. W. Stewart. Perturbation analyses for the QR factorization. SIAM Journal on Matrix Analysis and Applications, 18 (3):775–791, 1997. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

deBoor:1978:SRA

[dBG78] Carl de Boor and Gene H. Golub, editors. Symposium on Recent Advances in Numerical Analysis (1978: Madison, Wis.). Recent Advances in Numerical Analysis: Proceedings of a Symposium Conducted by the Mathematics Research Center, the University of Wisconsin, Madison, May 22– 24, 1978, volume 41 of Publication of the Mathematics Research Center, the University of Wisconsin, Madison. Army Mathematics Research Center, Academic Press Inc., New York, NY, USA, 1978. ISBN 0-12-208360-1. LCCN QA3 .U45 no. 41; QA297 S994 1978. URL http://

                                        catalog.hathitrust.org/api/
                                        volumes/oclc/65991765.html;
                                         http://www.gbv.de/dms/hbz/
                                        toc/ht001227727.pdf.

Dongarra:1979:LUG

[DBMS79] J. J. Dongarra, J. R. Bunch, C. B. Moler, and G. W. Stewart. LIN-PACK Users’ Guide. SIAM (Society for Industrial and Applied Mathematics), Philadelphia, PA, USA, 1979. ISBN 0-89871-172-X (paperback). 320 pp. LCCN QA76.73 .L22 L5 1979; QA184 .L56 1982; QA214 .L56 1979.

Daniel:1976:RSA

[DGKS76] J. W. Daniel, W. B. Gragg, L. Kaufman, and G. W. Stewart. Reorthogonalization and stable algorithms for updating the Gram–Schmidt QR factorization. Mathematics of Computation, 30(136):772–795, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dejon:1969:CAF

[DH69] Bruno Dejon and Peter Henrici, editors. Constructive Aspects of the Fundamental Theorem of Algebra: proceedings of a symposium conducted at the IBM Research Laboratory, Zürich-Rüschlikon, Switzerland, June 5– 7, 1967. Wiley, New York, NY, USA, 1969. ISBN 0-471-20300-9. LCCN QA212 .C65.

Duff:1979:SMP

[DS79] Iain S. Duff and G. W. Stewart, editors. Sparse Matrix. Proceedings 1978: Symposium held in the Hyatt Regency, Knoxville, Tennessee on November 2–3, 1978. SIAM (Society for Industrial and Applied Mathematics), Philadelphia, PA, USA, 1979. ISBN 0-89871-160-6. LCCN QA188 S9

1978. URL http://www.gbv.de/
dms/hbz/toc/ht000381636.pdf.

Dongarra:1984:LPS

[DS84] J. J. Dongarra and G. W. Stewart. LINPACK — A package for solving linear systems. In Cowell [Cow84], pages 20–48. ISBN 0-13-823501-5. LCCN QA76.95 .S68 1984.

Duff:2010:BGW

[Duf10]  Iain S. Duff.  Biography of

G. W. Stewart. In Kilmer and O’Leary [KO10], pages 3–9. ISBN 0-8176-4967-0 (hardcover), 0-8176-4968-9 (e-book). LCCN QA188 .S74 2010. URL http://

public.eblib.com/EBLPublic/
PublicView.do?ptiID=64586;
http://rave.ohiolink.edu/
ebooks/ebc/978081764968;
http://site.ebrary.com/id/
1042123.

Edwards:1988:CPC

[ER88] D. (David) Edwards and N. E.

(Niels E.) Raun, editors. COMP-STAT: proceedings in computational statistics, 8th symposium held in Copenhagen 1988. Physica-Verlag, Vienna, Austria; Heidelberg, Germany, 1988. ISBN 3-7908-0411-8. LCCN QA276.4 .C57

1988. URL http://catalog.
hathitrust.org/api/volumes/
oclc/19564603.html;   http:
//zbmath.org/?q=an:0732.00016.

Edelman:1993:SO

[ES93] Alan Edelman and G. W. Stewart. Scaling for orthogonality. IEEE Transactions on Signal Processing, 41(4):1676–1677, ???? 1993. CODEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=212741.

Fassbender:2006:CVF

[FGH^{+}06] Heike Fassbender, Michael Griebel, Olga Holtz, G. W. (Pete) Stewart, and Christoph Zenger. Curriculum vitae of Friedrich Ludwig Bauer. Linear Algebra and Its Applications, 417(2–3):299– 300, September 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special Issue in honor of Friedrich Ludwig Bauer.

Fraser:1995:BRBb

[Fra95]  Craig Fraser.  Book review:

Theoria combinationis observationum erroribus minimis obnoxiae [(Latin) [The theory of the combination of observations least subject to error]] by Carl Friedrich Gauss; G. W. Stewart. Isis, 86(4):660–661, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235423.

Flamm:1982:RHE

[FW82] David S. Flamm and Robert A.

Walker. Remark on “Algorithm 506: HQR3 and EXCHNG: Fortran subroutines for calculating and ordering the eigenvalues of a real upper Hessenberg matrix [F2]”. ACM Transactions on Mathematical Software, 8(2):219–220, June 1982. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Ste76a].

Griffin:1992:MNU

[GBS92] M. F. Griffin, E. C. Boman, and G. W. Stewart. Minimumnorm updating with the rankrevealing URV decomposition. In IEEE [IEE92], pages 293– 296 (vol. 5). ISBN 0-7803-0533-7 (hardcover), 0-7803-0532-9 (paperback), 0-7803-0534-5 (microfiche). LCCN TK 7882 S65 I16 1992. URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=226625.

Five volumes. IEEE catalog number 92CH3103-9.

Golub:1994:RAI

[GGL94] Gene Golub, Anne Greenbaum, and Mitchell Luskin, editors. Recent advances in iterative methods: [papers from the IMA Workshop on Iterative Methods for Sparse and Structured Problems, held in Minneapolis, Minnesota, February 24–March 1, 1992], volume 60 of The IMA Volumes in Mathematics and its Applications. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-94252-1 (New York), 3-540-94252-1 (Berlin). LCCN QA297.8 .R43

                                        1994. URL http://zbmath.org/
                                        ?q=an:0790.00015.

Golub:1987:GEY

[GHS87] G. H. Golub, Alan Hoffman, and G. W. Stewart. A generalization of the Eckart–Young–Mirsky matrix approximation theorem. Linear Algebra and Its Applications, 88/89(??):317–327, 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Golub:1988:GEY

[GHS88] G. H. Golub, A. Hoffman, and G. W. Stewart. A generalization of the Eckart–Young–Mirsky approximation theorem. Linear Algebra and Its Applications, 88/89: 317–328, 1988. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Golub:1976:RDLb

[GKS76a] G. H. Golub, Virginia Klema, and G. W. Stewart III. Rank degeneracy and least squares problems. Report TR-751, Department of Computer Science, University of Maryland, College Park, MD, USA, ???? 1976. ???? pp.

Golub:1976:RDLa

[GKS76b] Gene H. Golub, Virginia Klema, and G. W. Stewart. Rank degeneracy and least squares problems. Technical Report TR-456, University of Maryland, College Park, MD, USA, June 1976. ???? pp. Also published as Stanford University Computer Science Department Technical Report STAN-CS-76-559.

Gill:1985:PRB

[GMS^{+}85] P. E. Gill, W. Murray, M. A.

Saunders, G. W. Stewart, and M. H. Wright. Properties of a representation of a basis for the null space. Mathematical Programming, 33:172–186, 1985. CO-DEN MHPGA4. ISSN 0025-5610.

Guglielmi:2015:EAC

[GOS15] Nicola Guglielmi, Michael L.

Overton, and G. W. Stewart. An efficient algorithm for computing the generalized null space decomposition. SIAM Journal on Matrix Analysis and Applications, 36(1):38–54, ???? 2015. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Gragg:1976:SVS

[GS76] W. B. Gragg and G. W. Stewart. A stable variant of the secant method for solving nonlinear equations. SIAM Journal on Numerical Analysis, 13(6):889–903, December 1976. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Griffin:1991:UMR

[GS91] M. F. Griffin and G. W. Stewart. Updating MUSIC and root-MUSIC with the rankrevealing URV decomposition.

In Chen [Che91], pages 277– 281 (vol. 1). ISBN 0-8186-2470-1 (paperback), 0-8186-2471-X (microfiche), 0-8186-2472-8 (hardcover). LCCN TK 5102.5 A78 1991. URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=186456.

IEEE catalog number 91CH31120. IEEE Computer Society Press order number 2470.

Gauss:1995:TCO

[GS95] Carl Friedrich Gauss and G. W. (Gilbert W.) Stewart. Theory of the combination of observations least subject to error: part one, part two, supplement = Theoria combinationis observationum erroribus minimus obnoxiae: pars prior, pars posterior, supplementum, volume 11 of Classics in applied mathematics. SIAM (Society for Industrial and Applied Mathematics), Philadelphia, PA, USA, 1995. ISBN 0-89871-347-1 (paperback). xi + 241 pp. LCCN QA275 .G37313 1995. URL

http://www.loc.gov/catdir/
enhancements/fy0739/95006589-
d.html;   http://www.loc.gov/
catdir/enhancements/fy0739/
95006589-t.html. Translation

by G. W. Stewart of Gauss’ classic works.

George:1986:P

[GSV86a] Alan George, G. W. Stewart, and Robert Voigt. Preface. Linear Algebra and Its Applications, 77(??):1–2, May 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0024379586901588.

George:1986:PCS

[GSV86b] Alan George, G. W. (Gilbert W.)

Stewart, and Robert G. Voigt, editors. Parallel computing: special volume, volume 77 of Linear Algebra and Its Applications. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, May 1986. ISSN 0024-3795 (print), 1873-1856 (electronic). LCCN QA184 L75 v. 77.

Griffiths:1994:NAP

[GW94] D. F. (David Francis) Griffiths and G. A. (G. Alistair) Watson, editors. Numerical analysis 1993: proceedings of the 15th Dundee Conference, June–July 1993, volume 303 of Pitman research notes in mathematics series. Longman Scientific and Technical, Harlow, Essex, 1994. ISBN 0-582-22568-X. ISSN 0269-3674. LCCN QA297 .D85 1993.

Holland:2000:WLC

[HBC^{+}00] John H. Holland, Lashon B.

Booker, Marco Colombetti, Marco Dorigo, David E. Goldberg, Stephanie Forrest, Rick L. Riolo, Robert E. Smith, Pier Luca Lanzi, Wolfgang Stolzmann, and Stewart W. Wilson. What is a learning classifier system? Lecture Notes in Computer Science, 1813:3–??, 2000. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/

                                        0558/bibs/1813/18130003.htm;
                                         http://link.springer-ny.com/
                                        link/service/series/0558/papers/
                                        1813/18130003.pdf.

Huckaby:2003:CSQ

[HC03] David A. Huckaby and Tony F. Chan. On the convergence of Stewart’s QLP algorithm for approximating the SVD. Numerical Algorithms, 32(2–4):287–316, April 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

                                        ipsapp007.kluweronline.com/
                                        content/getfile/5058/45/10/
                                        abstract.htm;   http://ipsapp007.
                                        kluweronline.com/content/getfile/
                                        5058/45/10/fulltext.pdf.

Huckaby:2005:SPQ

[HC05] D. A. Huckaby and T. F. Chan. Stewart’s pivoted QLP decomposition for low-rank matrices. Numerical Linear Algebra with Applications, 12(2–3):153–159, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Householder:1969:CSI

[HS69a] A. S. Householder and G. W. Stewart. Comments on “Some iterations for factoring a polynomial”. Numerische Mathematik, 13(5):470–471, October 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Householder:1969:BHD

[HS69b] A. S. Householder and G. W. Stewart III. Bigradients, Hankel determinants, and the Padé table. In Dejon and Henrici [DH69], pages 131–150. ISBN 0-471-20300-9. LCCN QA212 .C65.

Householder:1971:NFP

[HS71] A. S. Householder and G. W.

Stewart. The numerical factorization of a polynomial. SIAM Review, 13(1):38–46, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hoberock:1972:IRP

[HS72] L. L. Hoberock and G. W. Stewart III. Input requirements and parametric errors for systems identification under periodic excitement. AMSE Transactions, 94 (??):296–302, ???? 1972. ISSN 0761-2532.

Higham:1987:NLA

[HS87] Nicholas J. Higham and G. W.

Stewart. Numerical linear algebra in statistical computing. In Iserles and Powell [IP87], pages 41–57. ISBN 0-19-853614-3. LCCN QA297.J65

1986. URL http://www.gbv.
de/dms/hbz/toc/ht002967923.
pdf;   http://zbmath.org/?
q=an:0611.00024.

Hoaglin:1976:CSS

[HW76] David C. (David Caster) Hoaglin and Roy E. (Roy Elmer) Welsch, editors. Proceedings of the Ninth Interface Symposium on Computer Science and Statistics / Harvard University, Massachusetts Institute of Technology, April 1–2, 1976. Prindle, Weber and Schmidt, Inc., Boston, MA, USA, 1976. ISBN 0-87150-237-2. LCCN QA276 C73 1976.

Iazeolla:1984:MCP

[ICH84] Giuseppe Iazeolla, Pierre Jacques Courtois, and A. Hordijk, editors. Mathematical computer performance and reliability: proceedings of the International workshop: Pisa, Italy, September 26– 30, 1983. North-Holland Publishing Co., Amsterdam, The Netherlands, 1984. ISBN 0-444-86892-5. LCCN QA76.9.E94 M37

1984. URL http://catalog.
hathitrust.org/api/volumes/
oclc/10505592.html;   http:
//www.dandelon.com/servlet/
download/attachments/dandelon/
ids/DE004D409A71E004A5EA4C12579DE00471484.
pdf.

IEEE:1991:ICA

[IEE91] IEEE, editor. 1991 International Conference on Acoustics, Speech, and Signal Processing: ICASSP-91: May 14–17, 1991, The Sheraton Centre Hotel and Towers, Toronto, Ontario, Canada. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-7803-0003-3 (paperback), 0-7803-0004-1 (hardcover), 0-7803-0005-X (microfiche). LCCN TA365 I58 1991; TK7882.S65 I16 1991. URL

http://ieeexplore.ieee.org/
servlet/opac?punumber=540.

IEEE:1992:III

[IEE92] IEEE, editor. ICASSP-92 / 1992 IEEE International Conference on Acoustics, Speech, and Signal Processing, March 23– 26, 1992, the San Francisco Marriott, San Francisco, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-7803-0533-7 (hardcover), 0-7803-0532-9 (paperback), 0-7803-0534-5 (microfiche). LCCN TK 7882 S65 I16 1992. Five volumes. IEEE catalog number 92CH3103-9.

IEEE:2002:IIC

[IEE02] IEEE, editor. 2002 IEEE international conference on acoustics, speech and signal processing: proceedings, May 13-17, 2002, Renaissance Orlando Resort, Orlando, Florida, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7803-7402-9 (paperback). LCCN TA365 I11 2002; TK7882.S65 I61 2002. IEEE catalog number 02CH37334C.

Iserles:1987:SAN

[IP87] A. Iserles and M. J. D. Powell, editors. State of the Art in Numerical Analysis. Proceedings of the Joint IMA/SIAM Conference held at the University of Birmingham, 14–18 April 1986, volume 9 of The Institute of Mathematics and Its Applications conference series; new series. Oxford University Press, Oxford, UK, 1987. ISBN 0-19-853614-3. LCCN QA297.J65

                                        1986. URL http://www.gbv.
                                        de/dms/hbz/toc/ht002967923.
                                        pdf;   http://zbmath.org/?
                                        q=an:0611.00024.

Jia:2001:ARR

[JS01] Zhongxiao Jia and G. W. Stewart. An analysis of the Rayleigh– Ritz method for approximating eigenspaces. Mathematics of Computation, 70(234):637–647, April 2001. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

                                        http://www.ams.org/journal-
                                        getitem?pii=S0025-5718-00-
                                        01208-4;   http://www.ams.
                                        org/mcom/2001-70-234/S0025-
                                        5718-00-01208-4/S0025-5718-
                                        00-01208-4.dvi;   http://www.
                                        ams.org/mcom/2001-70-234/S0025-
                                        5718-00-01208-4/S0025-5718-
                                        00-01208-4.pdf;   http://www.
                                        ams.org/mcom/2001-70-234/S0025-
                                        5718-00-01208-4/S0025-5718-
                                        00-01208-4.ps;   http://www.
                                        ams.org/mcom/2001-70-234/S0025-
                                        5718-00-01208-4/S0025-5718-
                                        00-01208-4.tex.

Kincaid:1997:BRA

[Kin97] David R. Kincaid. Book review: Afternotes on Numerical Analysis, by G. W. Stewart. SIAM Review, 39(1):153, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

King:1999:BRA

[Kin99] J. Thomas King. Book reviews: Afternotes goes to graduate school, lectures in advanced numerical analysis, by G. W. Stewart. Mathematics of Computation, 68(226):887–891, April 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-
01060-1&u=/mcom/1999-68-226/

.

Kilmer:2010:GWS

[KO10] Misha Elena Kilmer and Dianne P. O’Leary, editors. G. W. Stewart: selected works with commentaries. Contemporary mathematicians. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2010. ISBN 0-8176-4967-0 (hardcover), 0-8176-4968-9 (e-book). xii + 729 pp. LCCN QA188 .S74 2010. URL http://

public.eblib.com/EBLPublic/
PublicView.do?ptiID=64586;
 http://rave.ohiolink.edu/
ebooks/ebc/978081764968;
http://site.ebrary.com/id/
1042123.

Kaagstrom:1983:MPP

[KR83] Bo Kågström and Axel Ruhe, editors. Matrix Pencils: Proceedings of a Conference Held at Pite Havsbad, Sweden, March 22– 24, 1982, volume 973 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 973. URL http:/

                                        /link.springer.com/book/10.
                                        1007/BFb0062089;   http://
                                        www.springerlink.com/content/
                                        978-3-540-39447-1.

Kilmer:1999:IRM

[KS99] Misha Kilmer and G. W. Stewart. Iterative regularization and MINRES. SIAM Journal on Matrix Analysis and Applications, 21(2):613–628, 1999. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/34862.

Lam:2013:DFP

[LHS13] Michael O. Lam, Jeffrey K. Hollingsworth, and G. W. Stewart. Dynamic floating-point cancellation detection. Parallel Computing, 39(3):146–155, March 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0167819112000622.

Liu:1993:AEB

[LOSW93] K. J. R. Liu, D. P. O’Leary, G. W. Stewart, and Y.-J. J. Wu. An adaptive ESPRIT based on URV decomposition. In IEEE [IEE02], pages 37–40 (vol. 4). ISBN 0-7803-7402-9 (paperback). LCCN TA365 I11 2002; TK7882.S65 I61 2002. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=319588.

IEEE catalog number 02CH37334C.

Liu:1994:UET

[LOSW94] K. J. R. Liu, D. P. O’Leary, G. W. Stewart, and Y.-J. J. Wu. URV ESPRIT for tracking time-varying signals. IEEE Transactions on Signal Processing, 42(12):3441–3448, ???? 1994. CODEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=340778.

Latouche:1995:NMM

[LS95] G. Latouche and G. W. Stewart.

Numerical methods for M/G/1 type queues. In Stewart [Ste95g], pages 571–581. ISBN 0-7923-9550-6. LCCN QA274.7 .I595

1995. URL http://zbmath.org/
?q=an:0940.00042.

Li:2000:NRP

[LS00] Ren-Cang Li and G. W. Stewart.

A new relative perturbation theorem for singular subspaces. Linear Algebra and Its Applications, 313(1–3):41–51, July 1, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/30/19/130/25/
27/abstract.html;   http://
www.elsevier.nl/gej-ng/10/
30/19/130/25/27/article.pdf.

Lee:2009:AET

[LS09] Che-Rung Lee and G. W. Stewart. Algorithm 879: EIGEN-TEST — a test matrix generator for large-scale eigenproblems. ACM Transactions on Mathematical Software, 35(1):7:1–7:11, July 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Moonen:1993:LAL

[MGD93] Marc S. Moonen, Gene H. Golub, and Bart L. R. De Moor, editors. Linear Algebra for Large Scale and Real-Time Applications: Proceedings of the NATO Advanced Study Institute, Leuven, Belgium, August 3–14, 1992, volume 232 of NATO ASI series. Series E, Applied sciences. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993. ISBN 0-7923-2151-

0. LCCN QA185.D37 L56 1993.

URL http://catdir.loc.gov/
catdir/enhancements/fy0823/
92046135-d.html;   http://
www.gbv.de/dms/hbz/toc/ht004938330.
pdf;   http://zbmath.org/?
q=an:0810.00029.

Miller:1975:TNA

[Mil75] John J. H. Miller, editor. Topics in numerical analysis II: proceedings of the Royal Irish Academy Conference on Numerical Analysis, 1974. Published for the Royal Irish Academy by Academic Press, London, UK, 1975. ISBN 0-323-14134-X, 0-12-496952-6 (e-book), 0-323-14134-X (e-book). LCCN QA297 .R69. URL http://

catalog.hathitrust.org/api/
volumes/oclc/6100019.html;
 http://www.sciencedirect.
com/science/book/9780124969520.

Meyer:1993:LAM

[MP93] C. D. (Carl Dean) Meyer and Robert J. Plemmons, editors. Linear algebra, Markov chains, and queueing models: [… some of the lectures given at the Workshop Linear Algebra, Markov Chains, and Queueing Models held January 13–17, 1992, as part of the Year of Applied Linear Algebra at the Institute for Mathematics and its Applications], volume 48 of The IMA volumes in mathematics and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-94085-5 (New York), 3-540-94085-5 (Berlin). LCCN QA184 .L545

1993. URL http://catalog.
hathitrust.org/api/volumes/
oclc/28182134.html;   http:
//www.gbv.de/dms/ilmenau/toc/
126000387.PDF;   http://zbmath.
org/?q=an:0779.00020.

Moler:1971:AGM

[MS71] Cleve B. Moler and Gilbert W.

Stewart. An algorithm for the generalized matrix eigenvalue problem Ax = λBx. Technical Report STAN-CS-71-232, Department of Computer Science, Stanford University, Stanford, CA, 1971. 50 pp. Issued jointly as report CNA 32 by the Center for Numerical Analysis, the University of Texas at Austin.

Moler:1972:AGM

[MS72] C. B. Moler and G. W. Stewart. An algorithm for generalized matrix eigenvalue problems. Report TR-3, University of Michigan, Ann Arbor, MI, USA, February 1972. 50 pp. URL http://

                                        zbmath.org/?q=an:0225.65046.

Moler:1973:AGM

[MS73] C. B. Moler and G. W. Stewart. An algorithm for generalized matrix eigenvalue problems. SIAM Journal on Numerical Analysis, 10(2):241–256, April 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

                                        zbmath.org/?q=an:0225.65046.

Moler:1978:HFA

[MS78] Cleve B. Moler and G. W. Stewart. On the Householder–Fox algorithm for decomposing a projection. Journal of Computational Physics, 28(1):82–91, July 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0021999178900487.

Moler:1979:EMF

[MS79] Cleve B. Moler and G. W. (Gilbert W.) Stewart. An efficient matrix factorization for digital image processing. Report LA-7637-MS, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, January 1979. 15 pp.

Meyer:1988:DPE

[MS88] Carl D. Meyer and G. W. Stewart. Derivatives and perturbations of eigenvectors. SIAM Journal on Numerical Analysis, 25(3):

679–691, June 1988. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mathias:1993:BAS

[MS93a] R. Mathias and G. W. Stewart. A block QR algorithm and the singular value decomposition. Linear Algebra and Its Applications, 182:91–100, March 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mathias:1993:BQA

[MS93b] R. Mathias and G. W. Stewart. A block QR algorithm and the singular value decomposition. Linear Algebra and Its Applications, 188:91–100, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McAllister:1984:RRR

[MSS84] D. F. McAllister, G. W. Stewart, and W. J. Stewart. On a Rayleigh–Ritz refinement technique for nearly uncoupled stochastic matrices. Linear Algebra and Its Applications, 60 (??):1–25, August 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379584900673.

OLeary:1990:BSP

[O’L90] Dianne P. O’Leary. On bounds for scaled projections and pseudoinverses. Linear Algebra and Its Applications, 132:115–117, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See [Ste89].

OLeary:1986:MCM

[OPSW86] Dianne P. O’Leary, Roger Pierson, G. W. Stewart, and Mark Weiser. The Maryland Crab: A module for building parallel computers. Technical Report CS-1660, Department of Computer Science, University of Maryland, College Park, MD, USA, 1986. Also issued as University of Maryland Institute for Advanced Computer Studies Report UMIACS-86-9.

OLeary:1985:DFA

[OS85] Dianne P. O’Leary and G. W. Stewart. Data-flow algorithms for parallel matrix computation. Communications of the Association for Computing Machinery, 28(8):840–853, August 1985. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.
acm.org/pubs/toc/Abstracts/
0001-0782/4025.html.

OLeary:1986:ASP

[OS86] Dianne P. O’Leary and G. W. Stewart. Assignment and scheduling in parallel matrix factorization. Linear Algebra and Its Applications, 77(??):275–299, May 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0024379586901722.

Special volume on parallel computing.

OLeary:1987:DSA

[OS87] D. P. O’Leary and G. W.

Stewart. From determinacy to systaltic arrays. IEEE Transactions on Computers, C-36(11):1355–1359, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009475.

OLeary:1990:CEE

[OS90] D. P. O’Leary and G. W. Stewart. Computing the eigenvalues and eigenvectors of symmetric arrowhead matrices. Journal of Computational Physics, 90 (2):497–505, October 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0021999190901773.

Oleary:1998:CNR

[OS98] D. P. Oleary and G. W. Stewart. On the convergence of a new Rayleigh quotient method with applications to large eigenproblems. Electronic Transactions on Numerical Analysis (ETNA), 7:182–189, 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (elec-

tronic). URL http://etna.mcs.
kent.edu/vol.7.1998/pp182-
189.dir/pp182-189.pdf. Large

scale eigenvalue problems (Argonne, IL, 1997).

OLeary:1979:ELE

[OSV79] Dianne P. O’Leary, G. W. Stewart, and James S. Vandergraft. Estimating the largest eigenvalue of a positive definite matrix. Mathematics of Computation, 33 (148):1289–1292, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

OLeary:1986:DMP

[OSvdG86] Dianne P. O’Leary, G. W. Stewart, and Robert van de Geijn. DOMINO: a message passing environment for parallel computations. Technical Report TR-1648, Department of Computer Science, University of Maryland, College Park, MD, USA, April 1986.

OLeary:1989:DTO

[OSvdG89] Diane P. O’Leary, G. W. Stewart, and Robert van de Geijn. DOMINO: A transportable operating system for parallel computation. In Wouk [Wou89], pages 25–34. ISBN 0-89871-238-6. LCCN QA76.5 .W665 1986.

                                        URL http://books.google.
                                        com/books?id=O3GOAAAAIAAJ;
                                         http://catalog.hathitrust.
                                        org/api/volumes/oclc/19630560.
                                        html;   http://zbmath.org/?
                                        q=an:0681.68004.

Parlett:1981:BRBb

[Par81] B. N. Parlett. Book review: LIN-PACK Users’ Guide (J. J. Dongarra, J. R. Bunch, C. B. Moler and G. W. Stewart). SIAM Review, 23(1):126–128, ???? 1981.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Plemmons:1981:RBM

[Ple81] R. J. Plemmons. Review of Sparse matrix proceedings 1978, edited by Iain S. Duff and G. W. Stewart. Linear Algebra and Its Applications, 40(??):275–278, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0024379581901580.

Priest:2004:ESC

[Pri04] Douglas M. Priest. Efficient scaling for complex division. ACM Transactions on Mathematical Software, 30(4):389–401, December 2004. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Rice:1977:MSS

[Ric77] John R. (John Rischard) Rice, editor. Mathematical Software Symposium, University of Wisconsin, Madison, 1977. Mathematical Software III: Proceedings of a Symposium Conducted by the Mathematics Research Center, the University of Wisconsin, Madison, March 28–30, 1977. Number 39 in Publication of the Mathematics Research Center, the University of Wisconsin, Madison. Academic Press Inc., New York, NY, USA, 1977. ISBN 0-12-587260-7 (hardcover). ix + 388 pp. LCCN QA3 .U45 no. 39; QA297 .M36 1977. URL http://

                                        zbmath.org/?q=an:0234.68003.

Rosenfeld:1974:IPP

[Ros74] Jack L. Rosenfeld, editor. Information processing, 1974; Proceedings of IFIP Congress 74, Stockholm, Sweden, August 5–10, 1974. North-Holland Publishing Co., Amsterdam, The Netherlands, 1974. ISBN 0-444-10689-8, 0-7204-2803-3. LCCN QA 76 I615.

Roginsky:1998:TDE

[RS98] Jacob Roginsky and G. W. Stewart. Time-domain extraction of broad-band sources by Tikhonov– Phillips regularization of triangular Toeplitz kernels. Technical Report UMIACS-TR-95-87, Institute for Advanced Computer Studies, University of Maryland, College Park, MD, USA, October 15, 1998. URL http://drum.

                                        lib.umd.edu/handle/1903/755.

Serbin:1997:BRA

[Ser97] Steven M. Serbin. Book reviews: Afternotes on numerical analysis, by G. W. Stewart. Mathematics of Computation, 66 (219):1367–1374, July 1997. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sharda:1989:IRC

[SGW^{+}89] R. Sharda, B. L. Golden, E. Wasil, O. Balci, and G. W. Stewart, editors. Impacts of Recent Computer Advances on Operations Research, volume 9 of Publications in Operations Research. North-Holland Publishing Co., Amsterdam, The Netherlands, 1989. ISBN 0-444-01492-6. LCCN T57.6 .I4831 1989.

URL http://zbmath.org/?q=
an:0727.90047.

Stewart:1968:NAN

[SL68] G. W. Stewart III and Dale W.

Lick. Numerical analysis: Numerical solution of a thin plate heat transfer problem. Communications of the Association for Computing Machinery, 11(9): 639–641, September 1968. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schilders:2008:BMU

[SMC08] W. H. A. Schilders, P. B. L. Meijer, and E. Ciggaar. Behavioural modelling using the MOESP algorithm, dynamic neural networks and the Bartels–Stewart algorithm. Applied Numerical Mathematics: Transactions of IMACS, 58(12):1972–1993, December 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Soderstrom:1974:NPI

[SS74] Torsten Söderström and G. W.

Stewart. On the numerical properties of an iterative method for computing the Moore–Penrose generalized inverse. SIAM Journal on Numerical Analysis, 11(1): 61–74, March 1974. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stewart:1990:MPT

[SS90] Gilbert Wright (Gilbert W.)

Stewart and Ji-guang Sun. Matrix Perturbation Theory. Computer science and scientific computing. Academic Press Inc., New York, NY, USA, 1990. ISBN 0-12-670230-6 (hardcover). xv + 365 pp. LCCN QA871 .S775 1990. URL http://

catalog.hathitrust.org/api/
volumes/oclc/21227976.html;
 http://www.loc.gov/catdir/
description/els032/90033378.
htm;   http://www.loc.gov/
catdir/description/els032/
90033378.html;   http://www.
loc.gov/catdir/toc/els031/
90033378.htm;   http://www.
loc.gov/catdir/toc/els031/
90033378.html.

Schweitzer:1993:LEP

[SS93] Paul J. Schweitzer and G. W. Stewart. The Laurent expansion of pencils that are singular at the origin. Linear Algebra and Its Applications, 183:237– 254, April 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stewart:1998:HTS

[SS98] Michael Stewart and G. W. Stewart. On hyperbolic triangularization: Stability and pivoting. SIAM Journal on Matrix Analysis and Applications, 19(4): 847–860, October 1998. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31958.

Saltzman:1989:IIP

[SSM89] M. J. Saltzman, R. Subramaniam, and R. E. Marsten. Implementing an interior point LP algorithm on a supercomputer. In Sharda et al. [SGW^{+}89], pages 158–168. ISBN 0-444-01492-6. LCCN T57.6 .I4831

1989. URL http://zbmath.org/
?q=an:0727.90047.

Stewart:1994:TSI

[SSM94] G. W. Stewart, W. J. Stewart, and D. F. McAllister. A two-stage iteration for solving nearly completely decomposable Markov chains. In Golub et al. [GGL94], pages 201–216. ISBN 0-387-94252-1 (New York), 3-540-94252-1 (Berlin). LCCN QA297.8

.R43 1994. URL http://zbmath.
org/?q=an:0790.00015.

Stenbakken:1989:AGT

[SSS89] G. N. Stenbakken, T. M. Souders, and G. W. Stewart. Ambiguity groups and testability. IEEE Transactions on Instrumentation and Measurement, 38(5):941–947, ???? 1989. CODEN IEIMAO. ISSN 0018-9456 (print), 1557-9662 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=39034.

Stewart:1967:GTF

[Ste67a] G. W. Stewart III. A generalization of a theorem of Fan on Gershgorin disks. Numerische Mathematik, 10:162, 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://zbmath.org/?q=
an:0158.03504.  See origi-

nal paper by Fan at https://

doi.org/10.1215/S0012-7094-

                                        58-02538-9, https://zbmath.

org/?qan:0081.25202=.

Stewart:1967:MDM

[Ste67b] G. W. Stewart III. A modification of Davidon’s minimization method to accept difference approximations to derivatives. Journal of the Association for Computing Machinery, 14(1):72–83, January 1967. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stewart:1968:STNa

[Ste68] Gilbert Wright Stewart III. Some topics in numerical analysis. Report ORNL-4303, Oak Ridge National Laboratory, Knoxville, TN, USA, September 1968. vi

                                        + 71 pp. URL http://www.
                                        ornl.gov/info/reports/1968/
                                        3445605155079.pdf.

Stewart:1969:AOI

[Ste69a] G. W. Stewart. Accelerating the orthogonal iteration for the eigenvectors of a Hermitian matrix. Numerische Mathematik, 13 (4):362–376, August 1969. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

                                        tronic). URL http://zbmath.
                                        org/?q=an:0185.40203.

Stewart:1969:CGI

[Ste69b] G. W. Stewart. On the continuity of the generalized inverse. SIAM Journal on Applied Mathematics, 17(1):33–45, January 1969. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stewart:1969:SIF

[Ste69c] G. W. Stewart. Some iterations for factoring a polynomial. Numerische Mathematik, 13(5): 458–470, October 1969. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://zbmath.
org/?q=an:0181.17401.

Stewart:1969:LMF

[Ste69d] G. W. Stewart III. On Lehmer’s method for finding the zeros of a polynomial. Mathematics of Computation, 23(108):829– 835, s24–s30, October 1969. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigendum [Ste71d].

Stewart:1969:STF

[Ste69e] G. W. Stewart III. A set theoretic formulation of backward rounding error analysis. Computation Center Report TNN-92, University of Texas at Austin, Austin, TX, USA, ???? 1969. ???? pp.

Stewart:1969:STNb

[Ste69f] Gilbert Wright Stewart III. Some Topics in Numerical Analysis. Ph.D. dissertation, University of Tennessee, Knoxville, Knoxville, TN, USA, May

1969. URL http://genealogy.
math.ndsu.nodak.edu/id.php?
id=58767;   http://search.
proquest.com/docview/302349990;
http://www.ornl.gov/info/
reports/1968/3445605155079.
pdf.

Stewart:1970:AAE

[Ste70a] G. W. Stewart. ACM Algorithm 384: Eigenvalues and eigenvectors of a real symmetric matrix [F2]. Communications of the Association for Computing Machinery, 13(6):369–371, June 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See errata [Ste70e].

Stewart:1970:IOS

[Ste70b] G. W. Stewart. Incorporating origin shifts into the QR algorithm for symmetric tridiagonal matrices. Communications of the Association for Computing Machinery, 13(6):365–367, June 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://zbmath.
org/?q=an:0195.45001.

Stewart:1970:SIF

[Ste70c] G. W. Stewart. On Samelson’s iteration for factoring polynomials. Numerische Mathematik, 15 (4):306–314, August 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://zbmath.
org/?q=an:0231.65050.

Stewart:1970:CSS

[Ste70d] G. W. Stewart. On the convergence of Sebastião e Silva’s method for finding a zero of a polynomial. SIAM Review, 12 (3):458–460, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://zbmath.
org/?q=an:0198.20905.

Stewart:1970:RAE

[Ste70e] G. W. Stewart. Remark on “Algorithm 384: Eigenvalues and eigenvectors of a real symmetric matrix [F2]”. Communications of the Association for Computing Machinery, 13(12):750, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Ste70a].

Stewart:1971:EAA

[Ste71a] G. W. Stewart. Error analysis of the algorithm for shifting the zeros of a polynomial by synthetic division. Mathematics of Computation, 25(113): 135–139, 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1971:EBA

[Ste71b] G. W. Stewart. Error bounds for approximate invariant subspaces of closed linear operators. SIAM Journal on Numerical Analysis, 8(4):796–808, December 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

zbmath.org/?q=an:0232.47010.

Stewart:1971:COA

[Ste71c] G. W. Stewart. On a companion operator for analytic functions. Numerische Mathematik, 18(1):26–43, February 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://zbmath.
org/?q=an:0211.18805.

Stewart:1971:CLM

[Ste71d] G. W. Stewart III. Corrigendum: “On Lehmer’s Method for Finding the Zeros of a Polynomial”. Mathematics of Computation, 25 (113):203, January 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [Ste69d].

Stewart:1972:SEP

[Ste72] G. W. Stewart. On the sensitivity of the eigenvalue problem Ax = λBx. SIAM Journal on Numerical Analysis, 9(4):669–686, December 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

                                        zbmath.org/?q=an:0252.65026.

Stewart:1973:CDM

[Ste73a] G. W. Stewart. Conjugate direction methods for solving systems of linear equations. Numerische Mathematik, 21(4):285–297, August 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1973:EPB

[Ste73b] G. W. Stewart. Error and perturbation bounds for subspaces associated with certain eigenvalue problems. SIAM Review, 15 (4):727–764, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stewart:1973:SIF

[Ste73c] G. W. Stewart. Some iterations for factoring a polynomial. II: A generalization of the secant method. Numerische Mathematik, 22(1):33–36, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1973:IMC

[Ste73d] G. W. (Gilbert W.) Stewart. Introduction to Matrix Computations. Computer Science and Applied Mathematics, Editor: Werner Rheinboldt. Academic Press Inc., New York, NY, USA, 1973. ISBN 0-12-670350-7. xiii + 441 pp. LCCN QA188 .S71

1973. URL http://catalog.
hathitrust.org/api/volumes/
oclc/658076.html;   http://
catdir.loc.gov/catdir/description/
els031/72082636.html;   http:
//catdir.loc.gov/catdir/toc/
els031/72082636.html;   http:
//www.gbv.de/dms/hbz/toc/ht000832348.
pdf;   http://www.loc.gov/
catdir/description/els031/
72082636.html;   http://www.
loc.gov/catdir/toc/els031/
72082636.htmlQ.

Stewart:1974:CMI

[Ste74a] G. W. Stewart. The convergence of multipoint iterations to multiple zeros. SIAM Journal on Numerical Analysis, 11(6): 1105–1120, December 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stewart:1974:MPE

[Ste74b] G. W. Stewart. Modifying pivot elements in Gaussian elimination. Mathematics of Computation, 28 (126):537–542, April 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1974:NTL

[Ste74c] G. W. Stewart. The numerical treatment of large eigenvalue problems. In Rosenfeld [Ros74], pages 666–672. ISBN 0-444-10689-8, 0-7204-2803-3. LCCN QA 76 I615.

Stewart:1975:CMC

[Ste75a] G. W. Stewart. The convergence of the method of conjugate gradients at isolated extreme points in the spectrum. Numerische Mathematik, 24(2):85–93, April 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1975:GTG

[Ste75b] G. W. Stewart. Gershgorin theory for the generalized eigenvalue problem Ax = λBx. Mathematics of Computation, 29(130):600– 606, April 1975. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1975:IPT

[Ste75c] G. W. Stewart. An inverse perturbation theorem for the linear least squares problem. ACM SIGNUM Newsletter, 10(2–3): 39–40, November 1975. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Stewart:1975:MSI

[Ste75d] G. W. Stewart. Methods of simultaneous iteration for calculating eigenvectors of matrices. In Miller [Mil75], pages 185–196. ISBN 0-323-14134-X, 0-12-496952-6 (e-book), 0-323-14134-X (e-book). LCCN QA297 .R69. URL http://

catalog.hathitrust.org/api/
volumes/oclc/6100019.html;
 http://www.sciencedirect.
com/science/book/9780124969520.

Stewart:1976:AHE

[Ste76a] G. W. Stewart. Algorithm 506:

HQR3 and EXCHNG: Fortran subroutines for calculating and ordering the eigenvalues of a real upper Hessenberg matrix [F2]. ACM Transactions on Mathematical Software, 2(3):275–280, September 1976. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [FW82].

Stewart:1976:BTL

[Ste76b] G. W. Stewart. A bibliographical tour of the large sparse generalized eigenvalue problem. In Bunch and Rose [BR76], pages 113–130. ISBN 0-12-141050-1. LCCN QA188 .S989 1975.

Stewart:1976:DCR

[Ste76c]  G. W. Stewart.  Detecting

and correcting rank degeneracy. SIAM Review, 18(4):829, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stewart:1976:ESP

[Ste76d] G. W. Stewart. The economical storage of plane rotations. Numerische Mathematik, 25(2):

137–138, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1976:SIC

[Ste76e] G. W. Stewart. Simultaneous iteration for computing invariant subspaces of non-Hermitian matrices. Numerische Mathematik, 25(2):123–136, June 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1976:CVM

[Ste76f] G. W. Stewart III. Compilers, virtual memory, and matrix computations. In Hoaglin and Welsch [HW76], pages 85–88. ISBN 0-87150-237-2. LCCN QA276 C73 1976.

Stewart:1977:PPI

[Ste77a] G. W. Stewart. On the perturbation of pseudo-inverses, projections, and linear least squares problems. SIAM Review, 19 (4):634–662, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stewart:1977:PBF

[Ste77b] G. W. Stewart. Perturbation bounds for the QR factorization of a matrix. SIAM Journal on Numerical Analysis, 14(3):509– 518, June 1977. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stewart:1977:RDL

[Ste77c] G. W. Stewart. Research, development, and LINPACK. In Rice [Ric77], pages 1–14. ISBN 0-12-587260-7 (hardcover). LCCN QA3 .U45 no. 39; QA297 .M36

1977. URL http://zbmath.org/
?q=an:0234.68003.

Stewart:1977:SCE

[Ste77d] G. W. Stewart. Sensitivity coefficients for the effects of errors in the independent variables in a linear regression. Technical Report TR-571, Department of Computer Science, University of Maryland, College Park, MD, USA, 1977.

Stewart:1978:BRV

[Ste78a] G. W. Stewart. Book review: Vorlesungen über numerische Mathematik, by Heinz Rutishauser. Bulletin of the American Mathematical Society, 84(4):660–664, July 1978. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Stewart:1978:PTG

[Ste78b] G. W. Stewart. Perturbation theory for the generalized eigenvalue problem. In de Boor and Golub [dBG78], pages 193– 206. ISBN 0-12-208360-1. LCCN QA3 .U45 no. 41; QA297 S994 1978. URL http://

catalog.hathitrust.org/api/
volumes/oclc/65991765.html;
http://www.gbv.de/dms/hbz/
toc/ht001227727.pdf.

Stewart:1978:SFS

[Ste78c] G. W. Stewart. SRRIT — A FORTRAN subroutine to calculate the dominant invariant subspaces of a real matrix. Technical Report TRR-514, Department of Computer Science, University of Maryland, College Park, MD, USA, 1978.

Stewart:1979:ERE

[Ste79a] G. W. Stewart. The effects of rounding error on an algorithm for downdating a Cholesky factorization. Journal of the Institute of Mathematics and its Applications, 23(2):203–213, March 1979. CODEN JMTAA8. ISSN 0020-2932.

Stewart:1979:NPS

[Ste79b] G. W. Stewart. A note on the perturbation of singular values. Linear Algebra and Its Applications, 28:213–216, December 1979. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379579901344.

Stewart:1979:PBD

[Ste79c] G. W. Stewart. Perturbation bounds for the definite generalized eigenvalue problem. Linear Algebra and Its Applications, 23:69–85, February 1979. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379579900946.

Stewart:1979:PH

[Ste79d] G. W. Stewart. Preface: A. S. Householder. Linear Algebra and Its Applications, 28:1– 3 + 1, December 1979. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379579901125.

Stewart:1980:BMI

[Ste80a] G. W. Stewart. The behavior of a multiplicity independent rootfinding scheme in the presence of error. BIT (Nordisk tidskrift for informationsbehandling), 20 (4):526–528, December 1980. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=20&issue=4&spage=526.

Stewart:1980:EGR

[Ste80b] G. W. Stewart. The efficient generation of random orthogonal matrices with an application to condition estimators. SIAM Journal on Numerical Analysis, 17(3): 403–409 (loose microfiche supplement), June 1980. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stewart:1981:CDH

[Ste81a] G. W. Stewart. Constrained definite Hessians tend to be well conditioned. Mathematical Programming, 21:235–238, 1981. CODEN MHPGA4. ISSN 0025-5610.

Stewart:1981:MCH

[Ste81b] G. W. Stewart. Matrix calculations on hand-held calculators. ACM SIGNUM Newsletter, 16(1):10–13, March 1981. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Stewart:1981:IDN

[Ste81c] G. W. Stewart. On the implicit deflation of nearly singular systems of linear equations. SIAM Journal on Scientific and Statistical Computing, 2(2):136– 140, June 1981. CODEN SIJCD4. ISSN 0196-5204.

Stewart:1982:CDP

[Ste82] G. W. Stewart. Computing the CS decomposition of a partitioned orthonormal matrix. Numerische Mathematik, 40(3):297– 306, December 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1983:CEB

[Ste83a] G. W. Stewart. Computable error bounds for aggregated Markov chains. Journal of the Association for Computing Machinery, 30(2): 271–285, April 1983. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stewart:1983:MCG

[Ste83b] G. W. Stewart. A method for computing the generalized singular value decomposition. Lecture Notes in Mathematics, 973:207–220, 1983. CO-DEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/
10.1007/BFb0062104/.

Stewart:1984:ABS

[Ste84a] G. W. Stewart. On the asymptotic behavior of scaled singular value and QR decompositions. Mathematics of Computation, 43(168):483–489, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1984:IPN

[Ste84b] G. W. Stewart. On the invariance of perturbed null vectors under column scaling. Numerische Mathematik, 44(1):61–65, June 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1984:SNU

[Ste84c] G. W. Stewart. On the structure of nearly uncoupled Markov chains. In Iazeolla et al. [ICH84], pages 287–302. ISBN 0-444-86892-5. LCCN QA76.9.E94 M37

1984. URL http://catalog.
hathitrust.org/api/volumes/
oclc/10505592.html;   http:
//www.dandelon.com/servlet/
download/attachments/dandelon/
ids/DE004D409A71E004A5EA4C12579DE00471484.
pdf.

Stewart:1984:RD

[Ste84d] G. W. Stewart. Rank degeneracy. SIAM Journal on Scientific and Statistical Computing, 5(2): 403–413, June 1984. CODEN SI-JCD4. ISSN 0196-5204.

Stewart:1984:SOP

[Ste84e] G. W. Stewart. A second order perturbation expansion for small singular values. Linear Algebra and Its Applications, 56 (??):231–235, January 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379584901289.

Stewart:1985:JLA

[Ste85a] G. W. Stewart. A Jacobilike algorithm for computing the Schur decomposition of a non-Hermitian matrix. SIAM Journal on Scientific and Statistical Computing, 6(4):853–864, October 1985. CODEN SIJCD4. ISSN 0196-5204.

Stewart:1985:NCD

[Ste85b] G. W. Stewart. A note on complex division. ACM Transactions on Mathematical Software, 11(3):238–241, September 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://

www.acm.org/pubs/citations/
journals/toms/1985-11-3/p238-

stewart/. See corrigendum [Ste86c] and the faster and more robust algorithm in [Pri04].

Stewart:1986:CSR

[Ste86a] G. W. Stewart. Collinearity, scaling, and rounding error. In Allen [All86], pages 195–198. ISBN 0-444-70018-8 (paperback). LCCN QA276.4 .S95 1985. URL http://

catalog.hathitrust.org/api/
volumes/oclc/13455449.html;
 http://www.dtic.mil/dtic/
tr/fulltext/u2/a166019.pdf;

http://zbmath.org/?q=an:0647.
62008.

Stewart:1986:CPA

[Ste86b] G. W. Stewart. Communication in parallel algorithms: An example. In Boardman and Stefanski [BS86], pages 11–14. LCCN QA 276.4 S95 1986; TL695 N67

1992. URL http://zbmath.org/
?q=an:0638.62003.

Stewart:1986:CNC

[Ste86c] G. W. Stewart. Corrigendum:

“A note on complex division”. ACM Transactions on Mathematical Software, 12(3):285, September 1986. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Ste85b].

Stewart:1987:CLSa

[Ste87a] G. W. Stewart. Collinearity and least squares regression. Statistical Science, 2(1):68–84, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.
ss/1177013439.

Stewart:1987:CLSb

[Ste87b] G. W. Stewart. [Collinearity and least squares regression]: Rejoinder. Statistical Science, 2 (1):98–100, February 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.
org/euclid.ss/1177013444.

Stewart:1987:ISC

[Ste87c] G. W. Stewart. Invariant subspaces and capital punishment (a participatory paper). Technical Report CS-TR 1923, Department of Computer Science, University of Maryland, College Park, MD, USA, 1987. ???? pp.

Stewart:1987:PIA

[Ste87d] G. W. Stewart. A parallel implementation of the QR-algorithm. Parallel Computing, 5(1–2):187– 196, July 1987. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

Stewart:1987:RBC

[Ste87e] G. W. Stewart. Review of Matrix Computations, by Gene H. Golub and Charles F. Van Loan. Linear Algebra and Its Applications, 95 (??):211–215, October 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/002437958790036X.

Stewart:1988:CCR

[Ste88a] G. W. Stewart. A curiosity concerning the representation of integers in noninteger bases. Mathematics of Computation, 51(184): 755–756, October 1988. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://zbmath.
org/?q=an:0699.10020.

Stewart:1988:PLA

[Ste88b] G. W. Stewart. Parallel linear algebra in statistical computing. In Edwards and Raun [ER88], pages 3–14. ISBN 3-7908-0411-8. LCCN QA276.4 .C57

1988. URL http://catalog.
hathitrust.org/api/volumes/
oclc/19564603.html;   http:
//zbmath.org/?q=an:0732.00016.

Stewart:1989:SPP

[Ste89] G. W. Stewart. On scaled projections and pseudo-inverses. Linear Algebra and Its Applications, 112(??):189–193, January 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379589905946.

See answer [O’L90] to a question raised in this paper.

Stewart:1990:CMC

[Ste90a] G. W. Stewart. Communication and matrix computations on large message passing systems. Parallel Computing, 16(1):27–40, November 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Stewart:1990:ICC

[Ste90b] G. W. Stewart. Incremental condition calculation and column selection. Technical Report CS-TR 2495, Department of Computer Science, University of Maryland, College Park, MD, USA, 1990. ???? pp.

Stewart:1990:SNU

[Ste90c] G. W. Stewart. On the sensitivity of nearly uncoupled Markov chains. In Stewart [Ste90i], pages 105–119. ISBN ???? LCCN ????

Stewart:1990:PTL

[Ste90d] G. W. Stewart. Perturbation theory and least squares with errors in the variables. In Brown and Fuller [BF90], pages 171– 181. ISBN 0-8218-5117-9, 0-8218-7700-3 (e-book). ISSN 0271-4132 (print), 1098-3627 (electronic). LCCN QA275 .A47

                                        1989. URL http://www.ams.
                                        org/conm/112/.

Stewart:1990:PTS

[Ste90e] G. W. Stewart. Perturbation theory for the singular value decomposition. Technical Report CS-TR 2539, Department of Computer Science, University of Maryland, College Park, MD, USA, 1990. ???? pp.

Stewart:1990:SPT

[Ste90f] G. W. Stewart. Stochastic perturbation theory. SIAM Review, 32 (4):579–610, December 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stewart:1990:UAS

[Ste90g] G. W. Stewart. An updating algorithm for subspace tracking. Technical Report CS-TR 2494, Department of Computer Science, University of Maryland, College Park, MD, USA, 1990. ???? pp. revised January 1991.

Stewart:1990:IMS

[Ste90h] G. W. (Pete) Stewart. An iterative method for solving linear inequalities. In Cox and Hammarling [CH90], pages 241– 247. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00.

URL http://zbmath.org/?q=

an:0234.68003. Based on papers from a conference in honour of the late James Hardy Wilkinson (died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th– 10th July 1987.

Stewart:1990:NSM

[Ste90i] W. J. Stewart, editor. Numerical Solutions of Markov Chains. Marcel Dekker, Inc., New York, NY, USA, 1990. ISBN ???? LCCN ????

Stewart:1991:FFT

[Ste91a] G. W. Stewart. FTP — file transfer program. ACM SIGNUM Newsletter, 26(4):2–3, October 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Stewart:1991:FTR

[Ste91b] G. W. Stewart. FTPing technical reports. NA Digest, 91(39):??, 1991. Electronic mail magazine:

na.help@na-net.ornl.gov.

Stewart:1991:JGP

[Ste91c] G. W. Stewart. Jeep: A general purpose style file. TEX and TUG News, 0(0):3–4, ???? 1991. ISSN 1065-240X. URL http:

                                        //tug.ctan.org/tex-archive/
                                        digests/ttn/ttn0n0.tex.

Stewart:1991:LLS

[Ste91d] G. W. Stewart. Lanczos and linear systems. Technical Report CS-TR 2641, Department of Computer Science, University of Maryland, College Park, MD, USA, 1991. ???? pp.

Stewart:1991:ARR

[Ste91e] G. W. Stewart. On an algorithm for refining a rank-revealing URV factorization and a perturbation theorem for singular values. Technical Report CS-TR 2626, Department of Computer Science, University of Maryland, College Park, MD, USA, 1991. ???? pp.

Stewart:1991:PTS

[Ste91f] G. W. Stewart. Perturbation theory for the singular value decomposition. In Vaccaro [Vac91], pages 99–109. ISBN 0-444-88896-9. LCCN TK5102.5 .S93

                                        1991. URL http://catalog.
                                        hathitrust.org/api/volumes/
                                        oclc/22911621.html.

Stewart:1991:TSR

[Ste91g] G. W. Stewart. Two simple residual bounds for the eigenvalues of a Hermitian matrix. SIAM Journal on Matrix Analysis and Applications, 12(2):205– 208, April 1991. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Stewart:1992:EAQ

[Ste92a] G. W. Stewart. Error analysis of QR updating with exponential windowing. Mathematics of Computation, 59(199):135– 140, July 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1992:NGS

[Ste92b] G. W. Stewart. Note on a generalized Sylvester equation. IMA Preprint Series #985, Institute for Mathematics and its Applications, University of Minnesota, Minneapolis, MN, USA, May 1992. 3 pp.

Stewart:1992:UAS

[Ste92c] G. W. Stewart. An updating algorithm for subspace tracking. IEEE Transactions on Signal Processing, 40(6):1535–1541, ???? 1992. CODEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=139256.

Stewart:1993:ASH

[Ste93a] G. W. Stewart. Alston Scott Householder (1904–1993). SIAM News, 26(??):??, October 1993. ISSN 0036-1437. URL http:/

/archive.is/20130115033809/
http://sites.uclouvain.be/
HHXIX/AlstonHouseholder/#selection-
215.0-215.36.

Stewart:1993:BRB

[Ste93b] G. W. Stewart. Book review:

Topics in Matrix Analysis (Roger A. Horn and Charles R. Johnson). SIAM Review, 35(1): 168–169, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stewart:1993:DRP

[Ste93c] G. W. Stewart. Determining rank in the presence of error. In Moonen et al. [MGD93], pages 275–291. ISBN 0-7923-2151-

0. LCCN QA185.D37 L56 1993.

URL http://catdir.loc.gov/
catdir/enhancements/fy0823/
92046135-d.html;   http://
www.gbv.de/dms/hbz/toc/ht004938330.
pdf;   http://zbmath.org/?
q=an:0810.00029.

Stewart:1993:GEP

[Ste93d] G. W. Stewart. Gaussian elimination, perturbation theory, and Markov chains. In Meyer and Plemmons [MP93], pages 59–69. ISBN 0-387-94085-5 (New York), 3-540-94085-5 (Berlin). LCCN QA184 .L545

1993. URL http://catalog.
hathitrust.org/api/volumes/
oclc/28182134.html;   http:
//www.gbv.de/dms/ilmenau/toc/
126000387.PDF;   http://zbmath.
org/?q=an:0779.00020.

Stewart:1993:EHS

[Ste93e] G. W. Stewart. On the early history of the singular value decomposition. SIAM Review, 35(4): 551–566, December 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stewart:1993:PCF

[Ste93f] G. W. Stewart. On the perturbation of LU, Cholesky, and QR factorizations. SIAM Journal on Matrix Analysis and Applications, 14(4):1141–1145, October 1993. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Stewart:1993:PMC

[Ste93g] G. W. Stewart. On the perturbation of Markov chains with nearly transient states. Numerische Mathematik, 65(1):135–141, May 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1993:URR

[Ste93h] G. W. Stewart. Updating a rank-revealing ULV decomposition. SIAM Journal on Matrix Analysis and Applications, 14(2): 494–499, April 1993. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Stewart:1994:D

[Ste94a] G. W. Stewart. UT V decompositions. In Griffiths and Watson [GW94], pages 225–236. ISBN 0-582-22568-X. ISSN 0269-3674. LCCN QA297 .D85 1993.

Stewart:1994:GSG

[Ste94b] G. W. Stewart. Gauss, statistics, and Gaussian elimination. Technical Report TR-3307, Department of Computer Science, University of Maryland, College Park, MD, USA, August 1994. 14 pp.

Stewart:1994:LLS

[Ste94c] G. W. Stewart. Lanczos and linear systems. In Brown et al. [BCEP94], pages 135–139. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Stewart:1994:GQD

[Ste94d] G. W. Stewart. On graded QR decompositions of products of matrices. Technical Report TR-3263, Department of Computer Science, University of Maryland, College Park, MD, USA, April 1994. 14 pp.

Stewart:1994:MCS

[Ste94e] G. W. Stewart. On Markov chains with sluggish transients. Technical Report TR-3306, Department of Computer Science, University of Maryland, College Park, MD, USA, June 1994. 8 pp.

Stewart:1994:CMI

[Ste94f] G. W. Stewart. On the convergence of multipoint iterations. Numerische Mathematik, 68(1): 143–147, June 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/00211/
                                        bibs/4068001/40680143.htm;
                                         http://science.springer.de/
                                        nmee/bibs/4068001/40680143.
                                        htm.

Stewart:1994:SSU

[Ste94g] G. W. Stewart. On the stability of sequential updates and downdates. Technical Report TR-3238, Department of Computer Science, University of Maryland, College Park, MD, USA, March 1994. 17 pp.

Stewart:1994:PTR

[Ste94h] G. W. Stewart. Perturbation theory for rectangular matrix pencils. Linear Algebra and Its Applications, 208/209:297–301, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stewart:1994:UUD

[Ste94i]  G. W. Stewart.  Updating

URV decompositions in parallel. Parallel Computing, 20(2): 151–172, February 24, 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.elsevier.
com/cgi-bin/cas/tree/store/
parco/cas_sub/browse/browse.
cgi?year=1994&volume=20&issue=
2&aid=828.

Stewart:1995:GSG

[Ste95a] G. W. Stewart. Gauss, statistics, and Gaussian elimination. Journal of Computational and Graphical Statistics, 4(1):1–11, March 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (elec-

tronic).  URL http://www.
jstor.org/stable/1390624.

Stewart:1995:IMS

[Ste95b] G. W. Stewart. An iterative method for solving linear inequalities. Technical Report CS-TR-1833, Department of Computer Science, University of Maryland, College Park, MD, USA, February 6, 1995. URL http://drum.

                                        lib.umd.edu/handle/1903/355.

Stewart:1995:GQD

[Ste95c] G. W. Stewart. On graded QR decompositions of products of matrices. Electronic Transactions on Numerical Analysis (ETNA), 3:39–49, 1995. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

                                        URL http://etna.mcs.kent.
                                        edu/vol.3.1995/pp39-49.dir/
                                        pp39-49.pdf.

Stewart:1995:PSC

[Ste95d] G. W. Stewart. On the perturbation of Schur complement in positive semidefinite matrix. Technical Report TR-95-38, Institute for Advanced Computer Studies, University of Maryland, College Park, MD, USA, 1995.

Stewart:1995:SBH

[Ste95e] G. W. Stewart. On the solution of block Hessenberg systems. Numerical Linear Algebra with Applications, 2(3):287– 296, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:

                                        //onlinelibrary.wiley.com/
                                        doi/10.1002/nla.1680020309/
                                        abstract.

Stewart:1995:SSU

[Ste95f] G. W. Stewart. On the stability of sequential updates and downdates. IEEE Transactions on Signal Processing, 43(11):2642–2648, ???? 1995. CODEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=482114.

Stewart:1995:CMC

[Ste95g] William J. Stewart, editor. Computations with Markov chains: proceedings of the 2nd International Workshop on the Numerical Solution of Markov Chains. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-9550-6. LCCN QA274.7 .I595 1995.

URL http://zbmath.org/?q=
an:0940.00042.

Stewart:1996:ANA

[Ste96] G. W. (Gilbert W.) Stewart.

Afternotes on numerical analysis: a series of lectures on elementary numerical analysis presented at the University of Maryland at College Park and recorded after the fact. SIAM (Society for Industrial and Applied Mathematics), Philadelphia, PA, USA, 1996. ISBN 0-89871-362-5 (paperback). x + 200 pp. LCCN QA297 .S785 1996. URL

http://www.loc.gov/catdir/
enhancements/fy0726/95047768-
d.html;   http://www.loc.gov/
catdir/enhancements/fy0726/
95047768-t.html.

Stewart:1997:BRA

[Ste97a] G. W. Stewart. Book review:

Accuracy and Stability of Numerical Algorithms, by Nicholas J. Higham. SIAM Review, 39 (1):164–166, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stewart:1997:EVN

[Ste97b] G. W. Stewart. Errors in variables for numerical analysts. In van Huffel [vH97], pages 3–10. ISBN 0-89871-393-5. LCCN QA275

                                        .R4 1997. URL http://zbmath.
                                        org/?q=an:0861.00018. Pro-

ceedings of the Second International Workshop on Total Least Squares and Errors-in-Variables Modeling, Leuven, Belgium, August 21–24, 1996.

Stewart:1997:MCS

[Ste97c] G. W. Stewart. On Markov chains with sluggish transients. Stochastic Models, 13:85–95, 1997. CO-DEN CSSME8. ISSN 1532-6349 (print), 1532-4214 (electronic).

Stewart:1997:PCF

[Ste97d] G. W. Stewart. On the perturbation of LU and Cholesky factors. IMA Journal of Numerical Analysis, 17(1):1–6, January 1997. CO-DEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

                                        URL http://www3.oup.co.uk/
                                        imanum/hdb/Volume_17/Issue_
                                        01/170001.sgm.abs.html.

Stewart:1997:WML

[Ste97e] G. W. Stewart. On the weighting method for least squares problems with linear equality constraints. BIT Numerical Mathematics, 37(4):961–967, December 1997. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/
BIT/contents/bit37.html.

Stewart:1997:TMG

[Ste97f] G. W. Stewart. The triangular matrices of Gaussian elimination and related decompositions. IMA Journal of Numerical Analysis, 17(1):7–16, January 1997. CO-DEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

URL http://www3.oup.co.uk/
imanum/hdb/Volume_17/Issue_
01/170007.sgm.abs.html.

Stewart:1998:SC

[Ste98a] G. W. Stewart. On sublinear convergence. Technical report UMIACS-TR-95-92, Institute for Advanced Computer Studies, University of Maryland, College Park, MD, USA, 1998/10/15/

1998. URL http://drum.lib.
umd.edu/handle/1903/760.

Stewart:1998:AM

[Ste98b] G. W. Stewart. On the adjugate matrix. Linear Algebra and Its Applications, 283 (1–3):151–164, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/laa/sub/1998/283/1-3/
6217.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=283&issue=
1-3&aid=6217.

Stewart:1998:TMG

[Ste98c]  G. W. Stewart.  The tri-

angular matrices of Gaussian elimination and related decompositions. Technical Report UMIACS-TR-95-91, Institute for Advanced Computer Studies, University of Maryland, College Park, MD, USA, October 15,

1998. URL http://drum.lib.
umd.edu/handle/1903/759.

Stewart:1998:TAE

[Ste98d] G. W. Stewart. Two algorithms for the efficient computation of truncated pivoted QR approximations to a sparse matrix. Technical Report UMIACS-TR-98-12, Institute for Advanced Computer Studies, University of Maryland, College Park, MD, USA, October 15, 1998. URL http://drum.

lib.umd.edu/handle/1903/941.

Stewart:1998:TSR

[Ste98e] G. W. Stewart. Two simple residual bounds for the eigenvalues of Hermitian matrices. Technical Report UMIACS-TR-89-123, Institute for Advanced Computer Studies, University of Maryland, College Park, MD, USA, October 15, 1998. URL http://drum.

lib.umd.edu/handle/1903/545.

Stewart:1998:AGG

[Ste98f] G. W. (Gilbert W.) Stewart. Afternotes goes to graduate school: lectures on advanced numerical analysis. SIAM (Society for Industrial and Applied Mathematics), Philadelphia, PA, USA, 1998. ISBN 0-89871-404-4 (paperback). xii + 245 pp. LCCN QA297 .S784 1998. URL

http://www.loc.gov/catdir/
enhancements/fy0726/97044449-

d.html;   http://www.loc.gov/
catdir/enhancements/fy0726/
97044449-t.html. A series of lec-

tures on advanced numerical analysis presented at the University of Maryland at College Park and recorded after the fact.

Stewart:1998:MAV

[Ste98g] G. W. (Gilbert W.) Stewart. Matrix algorithms. Volume 1. Basic decompositions. SIAM (Society for Industrial and Applied Mathematics), Philadelphia, PA, USA, 1998. ISBN 0-89871-414-1. xix + 458 pp. LCCN QA188 .S714 1998. Basic decompositions.

Stewart:1999:BRN

[Ste99a] G. W. Stewart. Book reviews:

Numerical linear algebra, by Lloyd N. Trefethen and David Bau, III. Mathematics of Computation, 68(225):??, January 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-
00992-8&u=/mcom/1999-68-225/

.

Stewart:1999:FAE

[Ste99b] G. W. Stewart. Four algorithms for the efficient computation of truncated pivoted QR approximations to a sparse matrix. Numerische Mathematik, 83(2):313–323, August 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer-ny.com/link/
service/journals/00211/bibs/

9083002/90830313.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083002/90830313.pdf.

Stewart:1999:QAS

[Ste99c] G. W. Stewart. The QLP approximation to the singular value decomposition. SIAM Journal on Scientific Computing, 20(4): 1336–1348, July 1999. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31951.

Stewart:2000:BRA

[Ste00a] G. W. Stewart. Book review: ARPACK Users’ Guide, Solution of Large-Scale Eigenvalue Problems with Implicitly Restarted Arnoldi Methods. Mathematics of Computation, 69(231):1309–1311, July 2000. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http:

//www.ams.org/mcom/2000-69-
231/S0025-5718-00-01249-7/
bookrev-S0025-5718-00-01249-
7.html;   http://www.ams.org/
mcom/2000-69-231/S0025-5718-
00-01249-7/S0025-5718-00-01249-
7.dvi;   http://www.ams.org/
mcom/2000-69-231/S0025-5718-
00-01249-7/S0025-5718-00-01249-
7.pdf;   http://www.ams.org/
mcom/2000-69-231/S0025-5718-
00-01249-7/S0025-5718-00-01249-
7.ps;   http://www.ams.org/
mcom/2000-69-231/S0025-5718-
00-01249-7/S0025-5718-00-01249-
7.tex.

Stewart:2000:DAM

[Ste00b] G. W. Stewart. The decompositional approach to matrix computation. Computing in Science and Engineering, 2(1):50–59, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

dlib.computer.org/cs/books/
cs2000/pdf/c1050.pdf;   http:
//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=814658;
 http://www.computer.org/cse/
cs1999/c1050abs.htm.

Stewart:2001:GST

[Ste01a] G. W. Stewart. A generalization of Saad’s theorem on Rayleigh– Ritz approximations. Linear Algebra and Its Applications, 327 (1–3):115–119, April 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
elsevier.nl/gej%2Dng/10/30/
19/152/25/35/abstract.html;
 http://www.elsevier.nl/gej%
2Dng/10/30/19/152/25/35/article.
pdf.

Stewart:2001:KSA

[Ste01b] G. W. Stewart. A Krylov– Schur algorithm for large eigenproblems. SIAM Journal on Matrix Analysis and Applications, 23(3):601–614, 2001. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/37152.

Stewart:2001:MAV

[Ste01c] G. W. Stewart. Matrix Algorithms. Volume II: Eigensystems. SIAM (Society for Industrial and Applied Mathematics), Philadelphia, PA, USA, 2001. ISBN 0-89871-503-2 (paperback), 0-89871-805-8 (e-book). xix + 469 pp. LCCN QA188 .S714 1998 v.2.

Stewart:2001:EGM

[Ste01d] G. W. Stewart. On the eigensystems of graded matrices. Numerische Mathematik, 90(2):349–370, December 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

                                        /link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        1090002/10900349.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        1090002/10900349.pdf.

Stewart:2002:AKS

[Ste02a] G. W. Stewart. Addendum to “A Krylov–Schur Algorithm for Large Eigenproblems”. SIAM Journal on Matrix Analysis and Applications, 24(2):599– 601, 2002. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/40315.

Stewart:2002:ARQ

[Ste02b] G. W. Stewart. Adjusting the Rayleigh quotient in semiorthogonal Lanczos methods. SIAM Journal on Scientific Computing, 24(1):201–207, January 2002. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/38898.

Stewart:2002:BEB

[Ste02c] G. W. Stewart. Backward error bounds for approximate Krylov subspaces. Linear Algebra and Its Applications, 340(1–3):81– 86, January 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/30/
19/172/27/32/abstract.html.

Stewart:2003:BOF

[Ste03a] G. W. Stewart. Building an oldfashioned sparse solver. Technical Report TR-4527, Department of Computer Science, University of Maryland, College Park, MD, USA, 2003. Also issued as University of Maryland Institute for Advanced Computer Studies UMI-ACS report TR-2003-95.

Stewart:2003:MLD

[Ste03b] G. W. Stewart. Memory leaks in derived types revisited. ACM Fortran Forum, 22(3):25–27, December 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Stewart:2003:PMP

[Ste03c] G. W. Stewart. On the powers of a matrix with perturbations. Numerische Mathematik, 96(2): 363–376, December 2003. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:2004:ELP

[Ste04] G. W. Stewart. An Elsner-like perturbation theorem for generalized eigenvalues. Linear Algebra and Its Applications, 390(1):1–5, October 1, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stewart:2005:EAQ

[Ste05] G. W. Stewart. Error analysis of the quasi-Gram–Schmidt algorithm. SIAM Journal on Matrix Analysis and Applications, 27 (2):493–506, April 2005. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/60779.

Stewart:2006:NGH

[Ste06] G. W. Stewart. A note on generalized and hypergeneralized projectors. Linear Algebra and Its Applications, 412(2–3):408– 411, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stewart:2008:BGS

[Ste08] G. W. Stewart. Block Gram– Schmidt orthogonalization. SIAM Journal on Scientific Computing, 31(1):761–775, ???? 2008. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Stewart:2009:FMP

[Ste09a] G. W. Stewart. Flap: A Matlab package for adjustable precision floating-point arithmetic. Report, Department of Computer Science, University of Maryland, College Park, MD, USA, 2009.

URL http://www.cs.umd.edu/~
stewart/flap/flap.html.

Stewart:2009:UCC

[Ste09b] G. W. Stewart. An unreliable convergence criterion for Arnoldi’s method. Technical Report CMSC TR-4938, Department of Computer Science, University of Maryland, College Park, MD, USA, 2009.

Stewart:2010:DF

[Ste10a] G. W. Stewart. On downdating QR factorizations. Technical report, Institute for Advanced Computer Studies, University of Maryland, College Park, MD, USA, 2010. 1 pp.

Stewart:2010:SAM

[Ste10b] G. W. Stewart. On the semidefinite B-Arnoldi method. SIAM Journal on Matrix Analysis and Applications, 31(3):1458–1468, ???? 2010. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL http:/

/link.aip.org/link/?SML/31/
1458/1.

Stewart:2011:FHS

[Ste11a] G. W. Stewart. FREDHOLM, HILBERT, SCHMIDT: Three fundamental papers on integral equations. Translated with commentary by G. W. Stewart. December 15, 2011. URL

http://www.umiacs.umd.edu/~
stewart/FHS.pdf.

Stewart:2011:NAO

[Ste11b] G. W. Stewart. On the numerical analysis of oblique projectors. SIAM Journal on Matrix Analysis and Applications, 32 (1):309–348, ???? 2011. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (elec-

tronic). URL http://epubs.
siam.org/simax/resource/1/
sjmael/v32/i1/p309_s1;   http:
//link.aip.org/link/?SML/32/
309/1.

Stewart:2011:WTE

[Ste11c]  G. W. Stewart.  When is

twice enough? (the oblique case). In Householder Symposium XVIII on Numerical Linear Algebra, June 12–17, 2011. Granlibakken Conference Center & Lodge, Tahoe City, California, pages 220–221. ????, ????, 2011. URL http://

www.maths.manchester.ac.uk/
~higham/conferences/householder/
HH11_Abstracts.pdf.

Stewart:2012:SLB

[Ste12] G. W. Stewart. Smooth local bases for perturbed eigenspaces. Technical report, Institute for Advanced Computer Studies, University of Maryland, College Park, MD, USA, 2012. 8 pp.

Stewart:2016:CJE

[Ste16a] G. W. Stewart. Camille Jordan 1875: Essay on the geometry of n dimensions. 2016.

URL http://www.umiacs.umd.
edu/~stewart/Jord75.pdf.

Stewart:2016:CCR

[Ste16b] G. W. Stewart. A canonical CS representation of a pair of subspaces. SIAM Journal on Matrix Analysis and Applications, 37 (2):539–549, 2016. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Stewart:1991:EGM

[SZ91a] G. W. Stewart and G. Zhang.

Eigenvalues of graded matrices and the condition numbers of a multiple eigenvalue. Numerische Mathematik, 58(7):703– 712, February 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1991:DMS

[SZ91b] G. W. Stewart and G. Zhang.

On a direct method for the solution of nearly uncoupled Markov chains. Numerische Mathematik, 59(1):1–11, April 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tomayko:1988:AOT

[TTB^{+}88] James E. Tomayko, Henry S.

Tropp, W. Buchholz, Grace Murray Hopper, C. Strachey, Eric A. Weiss, Robert M. Stewart, and Bernard A. Galler. Anecdotes: Origin of the term bit; origin of the word byte; the first bug; whence the “bug”; an impossible program; reminiscences of Los Alamos; Los Alamos from below; Jonathan Swift’s computing invention; debugging. Annals of the History of Computing, 10 (4):336–348, October/December 1988. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1988/pdf/a4336.pdf;   http:
//www.computer.org/annals/
an1988/a4336abs.htm.

Trevathan:1984:DAN

[TTH^{+}84] Charles E. Trevathan, Thomas D. Taylor, Raymond G. Hartenstein, Ann. C. Merwarth, and William N. Stewart. Development and application of NASA’s first standard spacecraft computer. Communications of the Association for Computing Machinery, 27(9):902–913, 1984. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vaccaro:1991:SSP

[Vac91] Richard J. Vaccaro, editor. SVD and signal processing, II: algorithms, analysis, and applications. Elsevier, Amsterdam, The Netherlands, 1991. ISBN 0-444-88896-9. LCCN TK5102.5 .S93

1991. URL http://catalog.
hathitrust.org/api/volumes/
oclc/22911621.html.

vanHuffel:1997:RAT

[vH97] Sabine van Huffel, editor. Recent advances in total least squares techniques and errors-in-variables modeling. SIAM (Society for Industrial and Applied Mathematics), Philadelphia, PA, USA, 1997. ISBN 0-89871-393-5. LCCN QA275 .R4 1997. URL http://

zbmath.org/?q=an:0861.00018.

Proceedings of the Second International Workshop on Total Least Squares and Errors-in-Variables Modeling, Leuven, Belgium, August 21–24, 1996.

vonMatt:1994:RES

[vMS94] Urs von Matt and G. W. Stewart.

Rounding errors in solving block Hessenberg systems. Technical Report TR-3345, Department of Computer Science, University of Maryland, College Park, MD, USA, September 1994. ???? pp.

vonMatt:1996:RES

[vS96] Urs von Matt and G. W. Stewart. Rounding errors in solving block Hessenberg systems. Mathematics of Computation, 65(213): 115–135, January 1996. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-
00667-9&u=/mcom/1996-65-213/

.

Wouk:1989:PPM

[Wou89] Arthur Wouk, editor. Parallel processing and medium-scale multiprocessors [Proceedings of the workshop, held in Stanford, California, January 1986]. SIAM (Society for Industrial and Applied Mathematics), Philadelphia, PA, USA, 1989. ISBN 0-89871-238-6. LCCN QA76.5 .W665 1986.

URL http://books.google.
com/books?id=O3GOAAAAIAAJ;
 http://catalog.hathitrust.
org/api/volumes/oclc/19630560.

                                        html;   http://zbmath.org/?
                                        q=an:0681.68004.